### v1

# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

Abstract
Label distribution learning (LDL) provides a framework wherein a distribution over categories rather than a single category is predicted, with the aim of addressing ambiguity in labeled data. Existing research on LDL mainly focuses on the task of point estimation, i.e., pinpointing an optimal distribution in the probability simplex conditioned on the input sample. In this paper, we estimate a probability distribution of all possible label distributions over the simplex, by unleashing the expressive power of the recently introduced Squared Neural Family (SNEFY). With the modeled distribution, label distribution prediction can be achieved by performing the expectation operation to estimate the mean of the distribution of label distributions. Moreover, more information about the label distribution can be inferred, such as the prediction reliability and uncertainties. We conduct extensive experiments on the label distribution prediction task, showing that our distribution modeling based method can achieve very competitive label distribution prediction performance compared with the state-of-the-art baselines. Additional experiments on active learning and ensemble learning demonstrate that our probabilistic approach can effectively boost the performance in these settings, by accurately estimating the prediction reliability and uncertainties.
1 Introduction
Label distribution learning (LDL) is a technique which handles ambiguity in multi-class classification, by utilising simplex-valued rather than categorical-valued labels in training data. Unlike traditional multi-class and multi-label learning paradigms, which assign a deterministic label prediction to instances, LDL corresponds to the question “How well does each of the labels describe an instance?”, by using a discrete probability distribution to characterize the instance’s affinity to all candidate categories. For example, when we predict the functionality of a district in a city, we might predict a result such as: the district has 20% functionality for business, 40% functionality for entertainment, and 40% functionality for education.
Many LDL algorithms have been proposed to directly predict label distribution vectors from instance features, by adapting machine learning algorithms designed for “hard” label prediction to the “soft” label prediction setting. Though a discrete distribution among candidate labels is predicted, existing LDL algorithms still operate at the level of point estimation, i.e., they search for a single point on a probability simplex (the set of all possible label distributions) for each given instance. The point estimation paradigm is particularly susceptible to data uncertainty and inexact mappings between instances and labels, due to the inherent complexity of the data collection and generation processes. Therefore, modeling the probability distribution of label distribution vectors, i.e., the probability distribution supported on the probability simplex, is an important step towards trustworthy LDL. An additional bonus of the distribution modeling is the ability to quantify the prediction reliability and uncertainties, which not only facilitates reliable model deployment in real-world safety critical applications, but is also essential to various reliability/uncertainty-aware tasks, like pseudo labeling, active learning, and ensemble learning.
Contributions
In this paper, we propose a novel LDL framework, SNEFY-LDL, by unleashing the probability modeling power of the recently proposed Squared Neural Family (SNEFY) (Tsuchida et al., 2023), a new class of tractable probability models. By restricting the support set of SNEFY to a probability simplex, SNEFY-LDL provides an expressive conditional distribution modeling of the label distribution vector given the input sample. The conditional distribution model has a closed-form normalizing constant, guaranteeing computational tractability. In this way, model parameters can be learned efficiently by maximizing the conditional likelihoods of training samples with stochastic gradient descent. Through computing the expectation of the conditional distribution, we can obtain the closed-form distribution mean, as a way to summarize the fitted model and predict label distributions for unlabeled samples. However, this is only one way to use the fitted model and the probability density values can be used to evaluate the reliability of label distribution predictions.
We conduct extensive experiments on fourteen real-world datasets to evaluate the label distribution prediction performance of the proposed SNEFY-LDL model. Experimental results show that SNEFY-LDL can achieve very competitive label distribution prediction performance, compared with the state-of-the-art LDL baselines. In addition, label distribution active learning and ensemble learning experiments are carried out to verify the efficacy of SNEFY-LDL in quantifying prediction uncertainties. The max-entropy principle is used to achieve active learning with the estimated SNEFY-LDL entropy, i.e., select the most informative unlabeled samples with the largest entropy values, query their labels and augment training samples, to attain the largest performance gain of the re-trained LDL model. Experimental results show that the max-entropy principle achieves significantly better active learning performance than the representativeness based active learning baselines. The experiments on ensemble learning demonstrate that SNEFY-LDL gives a further usecase for the fitted probabilistic model, as it provides an intelligent mechanism for weighting base learners, significantly outperforming the uniform weighting strategy.
2 Related Work
LDL is first proposed by (Geng et al., 2013) to solve the facial age estimation problem with insufficient training samples. Since then, a series of LDL algorithms have been developed, which can be divided into three categories: Problem Transformation (PT), Algorithm Adaptation (AA) and Specialized Algorithms (SA) (Geng, 2016).
Problem Transformation
PT (Geng, 2016) transforms the LDL problem into the single-label classification problem, by decomposing each training sample assigned with a label distribution into a set of duplicate training samples. Each of them is assigned with a different label and accounts for a partial sample in proportion to the label probability value, and is then used to train the single-label classifiers. The label likelihoods predicted by the single-label classifiers are then aggregated to form the final prediction of label distributions. PT-Bayes (Geng, 2016) and PT-SVM (Geng, 2016; Geng and Hou, 2015) transform the LDL problem into the single-label multi-class classification problem and respectively employ Bayes and SVM as the single-label classifiers. DF-LDL (González et al., 2021a) decomposes the label distribution prediction task into a number of one-versus-one binary classification tasks, and fuses the binary classification likelihoods to form the final label distribution prediction.
Algorithm Adaptation
AA (Geng, 2016) adapts traditional single-label classification models into the LDL setting, by leveraging the models’ compatibility in outputting a soft label distribution vector. Derived from the K Nearest Neighbor (KNN) algorithm (Wu et al., 2008), AA-KNN (Geng, 2016) predicts samples’ label distributions by averaging the label distributions of their $k$ nearest neighbors in feature space. AA-BP (Geng, 2016) constructs a three-layer neural network and adopts the softmax function as the activation of the output layer, making the neural network naturally produce a label distribution for each example. The neural network is trained by minimizing the sum of squared errors between the model output label distributions and the ground-truth label distributions.
Specialized Algorithms
SA (Geng, 2016) designs algorithms from scratch to directly solve the LDL problem. SA-IIS (Geng et al., 2013) and SA-BFGS (Geng, 2016) use the maximum entropy model to parameterize label distributions. They are trained by minimizing the Kullback-Leibler (KL) divergence between the model output and ground-truth label distributions, where Improved Iterative Scaling (IIS) (Malouf, 2002) and BFGS (Nocedal and Wright, 1999) are respectively leveraged by SA-IIS and SA-BFGS as optimizers. CPNN (Geng et al., 2013) uses a neural network to parameterize the joint probability distribution between sample features and labels following Modha’s probability distribution formulation (Modha and Fainman, 1994). BCPNN (Yang et al., 2017) and ACPNN (Yang et al., 2017) then improve on CPNN through leveraging binary label encoding and augmenting training samples respectively. LDLF (Shen et al., 2017) employs differentiable decision trees (Kontschieder et al., 2015) to model label distributions and KL divergence is used to design the learning objective. LDL-SCL (Jia et al., 2021) forces the label distributions of samples located closely in feature space to be similar to each other. LDL-LRR (Jia et al., 2023b) and LDL-DPA (Jia et al., 2023a) maintain the relative importance ranking between different labels in label distribution modeling, by penalizing a label importance ranking loss in their learning objectives.
Extensions
In addition, LDL has been extended to other tasks, like label enhancement (recovering label distributions from the one-hot single-label assignments) (Xu et al., 2019b, 2020; Zheng et al., 2023), multi-class classification (Wang and Geng, 2019, 2021b, 2021a), learning with incomplete supervision (Xu and Zhou, 2017), oversampling (González et al., 2021b), ordinal LDL (Wen et al., 2023), semi-supervised learning (Xie et al., 2023), and label calibration (He et al., 2024). Furthermore, LDL has been applied to solve numerous real-world problems, including facial age estimation (Geng et al., 2013), facial emotion recognition (Chen et al., 2020), head pose estimation (Xu et al., 2019a), crowd opinion prediction (Geng and Hou, 2015), emphasis selection (Shirani et al., 2019), lesion counting (Wu et al., 2019), and urban function distribution (Huang et al., 2023). It is worth noting that there is a related topic in the classical statistics literature, termed compositional data analysis (Greenacre, 2021), with a broad range of applications including geochemistry (e.g. labels correspond to mineral compositions) and ecology (e.g. labels correspond to relative abundance of species).
However, existing LDL algorithms mainly fall into the regime of point estimation. They discover an optimal discrete label distribution in the probability simplex with regard to a predefined learning objective, and do not provide the information about how prominent the optimal label distribution is, compared with the remaining distribution candidates. In this paper, we aim to model the distribution of label distributions, with the expectation that we can provide a promising mechanism for LDL uncertainty quantification.
3 Problem Definition
Assume we are given a set of $N$ i.i.d training samples $\mathcal{X}=\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{N}\}$ with each sample $\bm{x}\in\mathcal{X}$ located in the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. In addition, each sample $\bm{x}\in\mathcal{X}$ is described by a $L$-dimensional label distribution vector $\bm{\ell}_{\bm{x}}\in\mathbb{R}^{L}$ that takes values in the $(L-1)$-simplex $\Delta^{L-1}$, corresponding to a set of $L$ given labels $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$. The $l$-th entry $\bm{\ell}_{\bm{x}}^{y_{l}}$ of the label distribution vector $\bm{\ell}_{\bm{x}}$ corresponds to the probability of sample $\bm{x}$ belonging to the $l$-th class $y_{l}$, i.e., $\bm{\ell}_{\bm{x}}^{y_{l}}=\mathrm{P}(y_{l}|\bm{x})\geq 0$, satisfying the constraint that $\sum_{l=1}^{L}\bm{\ell}_{\bm{x}}^{y_{l}}=1$.
With the label distribution observations of training samples, our objective is to model the probability distribution of the label distribution vector $\bm{\ell}\in\Delta^{L-1}$ conditioned on any input sample $\bm{x}\in\mathbb{R}^{d}$, $\mathrm{P}(d\bm{\ell}|\bm{x})$.
For any sample $\bm{x}\in\mathbb{R}^{d}$ with an unknown label distribution, one can associate to it a label distribution vector $\bm{\ell}^{*}_{\bm{x}}$ obtained as the mean of the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x})$:
$\bm{\ell}^{*}_{\bm{x}}=\int_{\Delta^{L-1}}\bm{\ell}\,\mathrm{P}(d\bm{\ell}|\bm{x}),$ (1)
but the fitted model $\mathrm{P}(d\bm{\ell}|\bm{x})$ can serve many other purposes.
4 Preliminaries on SNEFY
Given a measure space $(\Omega,\mathcal{F},\mu)$ with set $\Omega$, sigma algebra $\mathcal{F}$, and nonnegative measure $\mu$, SNEFY defines a probability distribution $\mathrm{P}$ on some support $\mathcal{Z}\in\mathcal{F}$ to be proportional to the evaluation of the squared 2-norm of a neural network $\bm{f}$:
$\mathrm{P}(d\bm{z};\bm{V},\bm{\Theta})\triangleq\frac{{\|\bm{f}(\bm{t}(\bm{z});\bm{V},\bm{\Theta})\|}^{2}_{2}\mu(d\bm{z})}{\int_{\mathcal{Z}}{\|\bm{f}(\bm{t}(\bm{z});\bm{V},\bm{\Theta})\|}^{2}_{2}\mu(d\bm{z})},$ (2)
$\bm{f}(\bm{t}(\bm{z});\bm{V},\bm{\Theta})=\bm{V}\sigma(\bm{W}\bm{t}(\bm{z})+\bm{b}),\;\bm{\Theta}=(\bm{W},\bm{b}),$
where $\bm{t}(\cdot):\mathcal{Z}\rightarrow\mathbb{R}^{D}$ is the sufficient statistic, $\sigma$ is the activation function, $\bm{W}\in\mathbb{R}^{n\times D}$ and $\bm{b}\in\mathbb{R}^{n}$ are respectively the weight matrix and bias vector at the hidden layer of neural network $\bm{f}$, and $\bm{V}\in\mathbb{R}^{m\times n}$ are $\bm{f}$’s readout parameters. $\bm{\Theta}=(\bm{W},\bm{b})\in\mathbb{R}^{n\times(D+1)}$ denotes the concatenation of $\bm{W}$ and $\bm{b}$ and its $i$th row is denoted as $\bm{\theta}_{i}=(\bm{w}_{i},b_{i})\in\mathbb{R}^{D+1}$.
The distribution $\mathrm{P}(d\bm{z};\bm{V},\bm{\Theta})$ in Eq. (2) admits a more concise formulation
$\mathrm{P}(d\bm{z};\bm{V},\bm{\Theta})$ $=\frac{\mathrm{Tr}[\bm{V}^{\top}\bm{V}\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{z})]}{\mathrm{Tr}[\bm{V}^{\top}\bm{V}\bm{K}_{\bm{\Theta}}]}\mu(d\bm{z}),$ (3)
$=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{z}))}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}})}\mu(d\bm{z}),$
where $\widetilde{\bm{K}}_{\bm{\Theta}}(\bm{z})\in\mathbb{R}^{n\times n}$ is a positive semidefinite (PSD) matrix, whose $ij$th element is a kernel function of $\bm{\theta}_{i}$ and $\bm{\theta}_{j}$:
$\tilde{\bm{k}}_{\sigma,\bm{t}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{z})=\sigma(\bm{w}_{i}^{\top}\bm{t}(\bm{z})+b_{i})\sigma(\bm{w}_{j}^{\top}\bm{t}(\bm{z})+b_{j}),$ (4)
while $\bm{K}_{\bm{\Theta}}$ is the elementwise integral of $\widetilde{\bm{K}}_{\bm{\Theta}}(\bm{z})$, preserving the PSD property, with its $ij$th entry formulated as another kernel function of $\bm{\theta}_{i}$ and $\bm{\theta}_{j}$:
$\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})=\int_{\mathcal{Z}}\tilde{\bm{k}}_{\sigma,\bm{t}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{z})\mu(d\bm{z}).$ (5)
Under varying choices of the activation function $\sigma$, sufficient statistic $\bm{t}$, and the base measure $\mu$, $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})$ is able to be computed in closed form (see Table 1 of (Tsuchida et al., 2023)) in $\mathcal{O}(D)$. This makes SNEFY a tractable probability distribution model, with great expressivity and computational efficiency.
SNEFY also enjoys a closed-form formulation for conditional distributions, under mild conditions.
Theorem 1.
(Tsuchida et al., 2023) Let $\mathtt{z}=(\mathtt{z}_{1},\mathtt{z}_{2})$ jointly follow a SNEFY distribution, with support set $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, sufficient statistic $\bm{t}$, activation function $\sigma$, base measure $\mu$, as well as parameters $\bm{V}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})$. Assume that $\mu(d\bm{z})=\mu_{1}(d\bm{z}_{1})\mu_{2}(d\bm{z}_{2})$ and $\bm{t}(\bm{z})=(\bm{t}_{1}(\bm{z}_{1}),\bm{t}_{2}(\bm{z}_{2}))$. Then the conditional distribution of $\mathtt{z}_{1}$ given $\mathtt{z}_{2}=\bm{z}_{2}$ is also a SNEFY distribution with support set $\mathcal{Z}_{1}$, sufficient statistic $\bm{t}_{1}$, activation function $\sigma$, base measure $\mu_{1}$, as well as parameters $\bm{V}$ and $\bm{\Theta}_{1|2}=(\bm{W}_{1},\bm{W}_{2}\bm{t}_{2}(\bm{z}_{2})+\bm{b})$.
5 SNEFY-LDL
SNEFY provides an effective way to model the conditional distribution of the label distribution vector $\bm{\ell}\in\Delta^{L-1}$. We can assume that the concatenation of label distribution vector $\bm{\ell}\in\Delta^{L-1}$ and its conditioning sample $\bm{x}\in\mathbb{R}^{d}$, $\bm{z}=(\bm{\ell},\bm{x})$, follows a joint SNEFY distribution, with support set $\mathcal{Z}=\Delta^{L-1}\times\mathbb{R}^{d}$, sufficient statistic $\bm{t}(\bm{z})=(\bm{t}_{1}(\bm{\ell}),\bm{t}_{2}(\bm{x})):\mathcal{Z}\rightarrow\mathbb{R}^{D_{1}+D_{2}}$ composed of $\bm{t}_{1}(\cdot):\Delta^{L-1}\rightarrow\mathbb{R}^{D_{1}}$ and $\bm{t}_{2}(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{D_{2}}$, activation function $\sigma$, base measure $\mu(\bm{z})=\mu_{1}(\bm{\ell})\mu_{2}(\bm{x})$, as well as parameters $\bm{V}\in\mathbb{R}^{m\times n}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})\in\mathbb{R}^{n\times(D_{1}+D_{2}+1)}$. Following Theorem 1, given sample $\bm{x}$, the conditional distribution of its label distribution vector $\bm{\ell}$ is a SNEFY distribution with support set $\Delta^{L-1}$, sufficient statistic $\bm{t}_{1}$, activation function $\sigma$, base measure $\mu_{1}$, as well as parameters $\bm{V}$ and $\bm{\Theta}_{1|2}=(\bm{W}_{1},\bm{W}_{2}\bm{t}_{2}(\bm{x})+\bm{b})$. The conditional distribution is
$\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})\triangleq\frac{{\|\bm{f}(\bm{t}_{1}(\bm{\ell}),\bm{t}_{2}(\bm{x});\bm{V},\bm{\Theta})\|}^{2}_{2}\mu_{1}(d\bm{\ell})}{\int_{\Delta^{L-1}}{\|\bm{f}(\bm{t}_{1}(\bm{\ell}),\bm{t}_{2}(\bm{x});\bm{V},\bm{\Theta})\|}^{2}_{2}\mu_{1}(d\bm{\ell})},$ (6)
$\bm{f}\big{(}\bm{t}_{1}(\bm{\ell}),\bm{t}_{2}(\bm{x});\bm{V},\bm{\Theta}\big{)}=\bm{V}\sigma\big{(}\bm{W}_{1}\bm{t}_{1}(\bm{\ell})+\bm{W}_{2}\bm{t}_{2}(\bm{x})+\bm{b}\big{)}.$
Following Eq. (3), the distribution can be reformulated as
$\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ $=\frac{\mathrm{Tr}[\bm{V}^{\top}\bm{V}\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x})]}{\mathrm{Tr}[\bm{V}^{\top}\bm{V}\bm{K}_{\bm{\Theta}}(\bm{x})]}\mu_{1}(d\bm{\ell}),$ (7)
$=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x}))}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}\mu_{1}(d\bm{\ell}),$
where $\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x})\in\mathbb{R}^{n\times n}$ is a PSD matrix, with its $ij$th element being a kernel function of $\bm{\theta}_{i}=(\bm{w}_{1i},\bm{w}_{2i},b_{i})\in\mathbb{R}^{D_{1}+D_{2}+1}$ and $\bm{\theta}_{j}=(\bm{w}_{1j},\bm{w}_{2j},b_{j})\in\mathbb{R}^{D_{1}+D_{2}+1}$:
$\tilde{\bm{k}}_{\sigma,\bm{t}_{1},\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\ell},\bm{x})=\sigma(\bm{w}_{1i}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+b_{i})\cdot\sigma(\bm{w}_{1j}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{j}),$ (8)
where $\bm{w}_{1i}\in\mathbb{R}^{D_{1}}$ and $\bm{w}_{2i}\in\mathbb{R}^{D_{2}}$ are respectively the $i$th row of matrices $\bm{W}_{1}$ and $\bm{W}_{2}$, and $b_{i}$ is the $i$th element of the bias vector $\bm{b}$. Then $\bm{K}_{\bm{\Theta}}(\bm{x})$ is the elementwise integral of $\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x})$, preserving the PSD property, with its $ij$th entry formulated as another kernel function of $\bm{\theta}_{i}$ and $\bm{\theta}_{j}$:
$\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=\int_{\Delta^{L-1}}\tilde{\bm{k}}_{\sigma,\bm{t}_{1},\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\ell},\bm{x})\mu_{1}(d\bm{\ell}).$ (9)
By choosing the activation function $\sigma$, sufficient statistic $\bm{t}_{1}$ and the base measure $\mu_{1}$ carefully, the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ admits a closed form, which guarantees that the conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ is tractable. In particular, we have the following theorem:
Theorem 2.
Let $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}}):\Delta^{L-1}\rightarrow\mathbb{R}^{L}$ by setting $D_{1}=L$, the activation function $\sigma$ be the exponential function $\exp$, the base measure $\mu_{1}(d\bm{\ell})=d\bm{\ell}$ be the Lebesgue measure. Under the condition that $\bm{W}_{1}>-1/2$ elementwise, the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ admits a closed form:
$\bm{k}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}},$ (10)
where $w_{1il}$ is the $il$-th element of matrix $\bm{W}_{1}$ and $\Gamma(\cdot)$ is the gamma function.
The proof is provided in the Appendix. With the closed-form kernel function in Eq. (10), we can construct the conditional SNEFY distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in the form of Eq. (7). This model provides us with the freedom to choose any sufficient statistic $\bm{t}_{2}(\cdot)$ that is used to transform the input sample $\bm{x}$ from the original $d$-dimensional space to the latent $D_{2}$-dimensional space. To capture the non-linearity between input samples and their label distributions, deep neural networks can be leveraged to construct $\bm{t}_{2}(\cdot)$. The input can also be extended beyond the vector-format samples to data with special structures, like images, texts and graphs, where we can respectively leverage Convolutional Neural Networks (CNNs) (Venkatesan and Li, 2017), Transformers (Vaswani et al., 2017) and Graph Neural Networks (GNNs) (Kipf and Welling, 2017) to construct $\bm{t}_{2}(\cdot)$ for end-to-end learning.
The conditional distribution formulation $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ also provides a closed form of mean, variance and co-variance of the label distribution probabilities conditioned on the input sample $\bm{x}$. About this, we have the following theorem:
Theorem 3.
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (7) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
$\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{F}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$ (11)
where $\circ$ denotes Hadamard product, and $\bm{F}^{y_{r}}$ is a $n\times n$ matrix, whose $ij$th entry is
$F_{ij}^{y_{r}}=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})}.$ (12)
The conditional variance of $\bm{\ell}^{y_{r}}$, $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$, is
$\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{G}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}\\ -\mathrm{E}^{2}[\bm{\ell}^{y_{r}}|\bm{x}],$ (13)
where $\bm{G}^{y_{r}}$ is a $n\times n$ matrix , with its $ij$th element being
$G^{y_{r}}_{ij}=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$ (14)
For two different labels $y_{r}$ and $y_{s}$, with $y_{r}\neq y_{s}$, the conditional covariance of $\bm{\ell}^{y_{r}}$ and $\bm{\ell}^{y_{s}}$, $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$, is
$\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{H}^{y_{r},y_{s}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]\cdot\mathrm{E}[\bm{\ell}^{y_{s}}|\bm{x}],$ (15)
where $\bm{H}^{y_{r},y_{s}}$ is a $n\times n$ matrix , with its $ij$th element being
$H^{y_{r},y_{s}}_{ij}=\frac{(1+w_{1ir}+w_{1jr})(1+w_{1is}+w_{1js})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$ (16)
The proof is provided in the Appendix. Given the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$, the mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ can be used to predict the unknown label distribution as the expectation over all values in the simplex. We can use the variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to quantify label distribution prediction uncertainties. We can also use $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ with $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to construct confidence intervals for label distribution predictions by applying Chebyshev’s inequality (Grimmett and Stirzaker, 2020). The covariance $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$ is helpful for us to understand the correlations between two different labels. More importantly, all the statistics are conditioned on the input sample $\bm{x}$, guiding us to make instance-wise inferences.
The conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ relies on the parameters $\bm{V}$ and $\bm{\Theta}$, as well as the neural network parameters for constructing $\bm{t}_{2}$ (we also use $\bm{t}_{2}$ to denote the parameters without confusion). We train the model with maximum likelihood estimation (MLE), by minimizing the negative log conditional likelihoods on training samples:
$\min_{\bm{V},\bm{\Theta},\bm{t}_{2}}-\sum_{\bm{x}^{\prime}\in\mathcal{X}}\log\frac{\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})}{d\bm{\ell}}\bigg{|}_{\bm{x}=\bm{x}^{\prime},\bm{\ell}=\bm{\ell}_{\bm{x}^{\prime}}}.$ (17)
There are numerous metrics to measure the consistency between two label distributions, like Chebyshev distance, Kullback-Leibler divergence and Cosine coefficient (Geng, 2016). Instead of optimizing these metrics, the MLE based learning objective in Eq. (17) provides an alternative way to train the LDL model. The fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ can be applied to various downstream tasks, and its closed-form mean is able to offer accurate label distribution predictions that are generic to varying evaluation metrics.
Algorithm Description and Time Complexity
We train the SNEFY-LDL model with stochastic gradient descent. The training procedure is shown in Algorithm 1. The model parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ are first initialized by random numbers. We then iteratively select a batch of training samples, calculate the batched likelihoods with Eq. (7), and update parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ by descending them along the gradient of batched negative log likelihoods. Taking the epoch number as a constant and assuming the latent layers of $\bm{t}_{2}$ have neurons in the same scale as the neuron number in the last layer $D_{2}$, the time complexity of Algorithm 1 is $\mathcal{O}(N(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2}))$, which is linear to the number of training samples $N$, making the algorithm able to scale to large datasets. For any sample $\bm{x}$ with an unknown label distribution, its label distribution can be predicted fast in time complexity $\mathcal{O}(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2})$ using the closed-form mean formulation in Eq. (11).
Algorithm 1 Training SNEFY-LDL
Input: Training set $\{(\bm{x}_{1},\bm{\ell}_{\bm{x}_{1}}),(\bm{x}_{2},\bm{\ell}_{\bm{x}_{2}}),\cdots,(\bm{x}_{N},\bm{\ell}_{\bm{x}_{N}})\}$.
Parameter: $(\bm{V},\bm{\Theta},\bm{t}_{2})$.
Output: Optimized $(\bm{V},\bm{\Theta},\bm{t}_{2})$.
1:  $(\bm{V},\bm{\Theta},\bm{t}_{2})$leftarrow random initialization;
2:  while epoch number does not expire do
3:     $\mathcal{B}\leftarrow$ randomly split training set into batches;
4:     for each batch in $\mathcal{B}$ do
5:        Calculate batched $\bm{K}_{\bm{\Theta}}(\bm{x})$ with Eq. (10);
6:        Calculate batched likelihoods with Eq. (7);
7:        $(\bm{V},\bm{\Theta},\bm{t}_{2})$leftarrow update by descending along the gradient of batched negative log likelihoods;
8:     end for
9:  end while
10:  return optimized $(\bm{V},\bm{\Theta},\bm{t}_{2})$.
6 Experiments
In this section, extensive experiments are carried out to evaluate SNEFY-LDL’s performance in label distribution prediction and prediction uncertainty quantification.
Benchmark Datasets
We use fourteen datasets (Geng, 2016) to benchmark our experiments, including the Movie dataset containing label distributions on five movie rating scales (from one to five), the Natural Scene dataset with label distributions constructed by inconsistent multi-label ranking on natural scene images, the facial expression datasets SBU_3DFE and SJAFFE with label distributions on six emotions, as well as the ten Yeast series datasets built from ten biological experiments on a budding yeast, where labels are defined by different time points and label distributions are constructed by different gene expression levels at varying time points. The statistics of the fourteen benchmark datasets are summarized in Table 1.
Baseline Methods
We compare the proposed SNEFY-LDL model with eight state-of-the-art LDL baselines:
  • PT-SVM (Geng, 2016) transforms LDL into a multi-class classification problem and uses SVM as the classifier.
  • AA-KNN (Geng, 2016) adapts KNN from the single-label classification setting to the LDL setting.
  • CPNN (Geng et al., 2013) uses a neural network to model the feature-label joint distribution.
  • SA-BFGS (Geng, 2016) applies the maximum entropy model to characterize label distributions.
  • LDLF (Shen et al., 2017) employs the differentiable decision trees to model label distributions.
  • DF-LDL (González et al., 2021a) learns label distributions in a decomposition-fusion manner.
  • LDL-SCL (Jia et al., 2021) considers the label distribution correlations between similar samples.
  • LDL-LRR (Jia et al., 2023a) preserves the label importance ranking in label distribution modeling.
No. Dataset #Examples #Features #Labels
1 Movie 7,755 1,869 5
2 Natural Scene 2,000 294 9
3 SBU_3DFE 2,500 243 6
4 SJAFFE 213 243 6
5 Yeast_alpha 2,465 24 18
6 Yeast_cdc 2,465 24 15
7 Yeast_cold 2,465 24 14
8 Yeast_diau 2,465 24 7
9 Yeast_dtt 2,465 24 6
10 Yeast_elu 2,465 24 6
11 Yeast_heat 2,465 24 4
12 Yeast_spo 2,465 24 4
13 Yeast_spo5 2,465 24 3
14 Yeast_spoem 2,465 24 2
Table 1: Summary of the fourteen benchmark datasets.
Experimental Settings
Following (Geng, 2016), we use ten-fold cross-validation to evaluate the label distribution prediction performance, by randomly splitting each dataset into ten chunks, and iteratively choosing a chunk as the test set and taking the remaining nine chunks as the training set. LDL models are then trained on the training set and evaluated on the test set for ten rounds. Averaged performance is used to compare SNEFY-LDL with baseline methods.
The closed-form mean in Eq. (11) is used to make comparisons on the label distribution prediction task with the baselines which all return point estimates on the simplex. However, we note that SNEFY-LDL is a full conditional density model and its utility is broader than the baselines, as we illustrate in the active learning and ensemble learning experiments. When implementing SNEFY-LDL, $m$ and $n$ are respectively set to 32 and 64, $D_{2}$ is set as equal to $n$ and a one-layer neural network with ReLU activation is used to construct $\bm{t}_{2}$. The model is trained for 100 epochs with batch size 16. Weight clipping (Arjovsky et al., 2017) is used to control $\bm{W}_{1}>-1/2$ after each parameter update. For the eight baseline models, we use the implementations in the PyLDL library [^1] https://github.com/SpriteMisaka/PyLDL and keep the default hyperparameters.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.9332 $\bullet$ 0.9279 $\bullet$ 0.9244 $\bullet$ 0.9339 $\bullet$ 0.8596 $\bullet$ 0.9357 $\spadesuit$ 0.9036 $\bullet$ 0.9353 $\spadesuit$ 0.9354 -0.0002
Natural_Scene 0.7404 $\spadesuit$ 0.7112 $\bullet$ 0.6249 $\bullet$ 0.7168 $\bullet$ 0.7304 $\spadesuit$ 0.7377 $\spadesuit$ 0.6716 $\bullet$ 0.7455 $\spadesuit$ 0.7524 +0.0068
SBU_3DFE 0.9244 $\bullet$ 0.9215 $\bullet$ 0.9169 $\bullet$ 0.9383 $\bullet$ 0.9359 $\bullet$ 0.9386 $\bullet$ 0.9249 $\bullet$ 0.9284 $\bullet$ 0.9427 +0.0042
SJAFFE 0.9336 $\bullet$ 0.9490 $\spadesuit$ 0.9302 $\bullet$ 0.9543 $\spadesuit$ 0.9593 $\circ$ 0.9646 $\circ$ 0.9604 $\circ$ 0.9635 $\circ$ 0.9507 -0.0140
Yeast_alpha 0.9941 $\bullet$ 0.9938 $\bullet$ 0.9940 $\bullet$ 0.9946 $\circ$ 0.9943 $\spadesuit$ 0.9946 $\circ$ 0.9945 $\circ$ 0.9946 $\circ$ 0.9943 -0.0004
Yeast_cdc 0.9928 $\bullet$ 0.9924 $\bullet$ 0.9927 $\bullet$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9929 -0.0004
Yeast_cold 0.9879 $\bullet$ 0.9872 $\bullet$ 0.9875 $\bullet$ 0.9885 $\circ$ 0.9876 $\bullet$ 0.9884 $\spadesuit$ 0.9871 $\bullet$ 0.9885 $\circ$ 0.9883 -0.0002
Yeast_diau 0.9870 $\bullet$ 0.9867 $\bullet$ 0.9860 $\bullet$ 0.9879 $\circ$ 0.9873 $\spadesuit$ 0.9879 $\circ$ 0.9849 $\bullet$ 0.9879 $\circ$ 0.9873 -0.0006
Yeast_dtt 0.9934 $\bullet$ 0.9934 $\bullet$ 0.9938 $\spadesuit$ 0.9941 $\circ$ 0.9936 $\spadesuit$ 0.9939 $\spadesuit$ 0.9931 $\spadesuit$ 0.9941 $\circ$ 0.9939 -0.0002
Yeast_elu 0.9933 $\bullet$ 0.9932 $\bullet$ 0.9936 $\spadesuit$ 0.9940 $\circ$ 0.9939 $\circ$ 0.9941 $\circ$ 0.9939 $\circ$ 0.9940 $\circ$ 0.9936 -0.0005
Yeast_heat 0.9871 $\bullet$ 0.9867 $\bullet$ 0.9874 $\bullet$ 0.9880 $\circ$ 0.9877 $\spadesuit$ 0.9881 $\circ$ 0.9873 $\spadesuit$ 0.9880 $\circ$ 0.9878 -0.0003
Yeast_spo 0.9765 $\spadesuit$ 0.9728 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9746 $\bullet$ 0.9756 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9766 -0.0004
Yeast_spo5 0.9736 $\bullet$ 0.9711 $\bullet$ 0.9737 $\bullet$ 0.9742 $\spadesuit$ 0.9690 $\bullet$ 0.9698 $\bullet$ 0.9640 $\bullet$ 0.9742 $\spadesuit$ 0.9744 +0.0002
Yeast_spoem 0.9783 $\bullet$ 0.9764 $\bullet$ 0.9789 $\spadesuit$ 0.9790 $\spadesuit$ 0.9750 $\bullet$ 0.9694 $\bullet$ 0.8925 $\bullet$ 0.9790 $\spadesuit$ 0.9792 +0.0001
Table 2: The comparison of label distribution prediction performance measured by Cosine coefficient $\uparrow$.
Evaluation Metrics
Following (Geng, 2016), we adopt six metrics to evaluate label distribution prediction performance: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter). By denoting $\bm{\ell}$ and $\hat{\bm{\ell}}$ as ground-truth and predicted label distribution vectors respectively, these metrics are defined as follows:
  • $\mathrm{Cheby}(\bm{\ell},\hat{\bm{\ell}})=\|\bm{\ell}-\hat{\bm{\ell}}\|_{\infty}\downarrow$.
  • $\mathrm{Clark}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{2}\downarrow$.
  • $\mathrm{Canb}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{1}\downarrow$.
  • $\mathrm{KL}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\bm{\ell}^{y}\log\frac{\bm{\ell}^{y}}{\hat{\bm{\ell}}^{y}}\downarrow.$
  • $\mathrm{Cos}(\bm{\ell},\hat{\bm{\ell}})=\frac{\bm{\ell}^{\top}\hat{\bm{\ell}}}{\|\bm{\ell}\|_{2}\|\hat{\bm{\ell}}\|_{2}}\uparrow.$
  • $\mathrm{Inter}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\min(\bm{\ell}^{y},\hat{\bm{\ell}}^{y})\uparrow.$
For each metric, $\uparrow$ ($\downarrow$) indicates that higher (lower) scores imply better label distribution prediction performance.
Label Distribution Prediction
Table 2 shows the label distribution prediction performance measured by Cosine coefficient, as the averaged scores on ten-fold cross-validation. On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline. We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level. The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline.
From Table 2, we can find that the proposed SNEFY-LDL achieves very competitive label distribution prediction performance, compared with the state-of-the-art baselines. In the majority of cases, the performance of SNEFY-LDL is no worse than the compared baselines. On four datasets, SNEFY-LDL achieves the best Cosine coefficient scores. While SNEFY-LDL does not perform best on the remaining datasets, its performance gap with regard to the best performer is very marginal and even negligible. The performance scores measured by the remaining five metrics are provided in the Appendix, which show the same trend. The competitive label distribution prediction performance implies that the proposed SNEFY-LDL model is able to accurately model the distribution of label distribution vectors, offering an effective tool for label distribution prediction.
Method Cheby $\downarrow$ Clark $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
Random 0.1490$\pm$0.0098 $\bullet$ 0.6359$\pm$0.0247 $\bullet$ 1.1999$\pm$0.0448 $\bullet$ 0.1409$\pm$0.0136 $\bullet$ 0.9054$\pm$0.0093 $\bullet$ 0.7950$\pm$0.0089 $\bullet$
Kmeans 0.1456$\pm$0.0076 $\bullet$ 0.6242$\pm$0.0176 $\bullet$ 1.1815$\pm$0.0320 $\bullet$ 0.1361$\pm$0.0072 $\bullet$ 0.9079$\pm$0.0068 $\bullet$ 0.7981$\pm$0.0077 $\bullet$
CoreSet 0.1484$\pm$0.0164 $\bullet$ 0.6322$\pm$0.0383 $\bullet$ 1.1962$\pm$0.0720 $\bullet$ 0.1399$\pm$0.0179 $\bullet$ 0.9042$\pm$0.0158 $\bullet$ 0.7941$\pm$0.0171 $\bullet$
Graph Density 0.1428$\pm$0.0059 $\bullet$ 0.6175$\pm$0.0197 $\bullet$ 1.1671$\pm$0.0346 $\bullet$ 0.1326$\pm$0.0068 $\bullet$ 0.9108$\pm$0.0057 $\bullet$ 0.8013$\pm$0.0076 $\bullet$
Dirichlet 0.1472$\pm$0.0090 $\bullet$ 0.6282$\pm$0.0245 $\bullet$ 1.1899$\pm$0.0470 $\bullet$ 0.1376$\pm$0.0116 $\bullet$ 0.9064$\pm$0.0085 $\bullet$ 0.7959$\pm$0.0099 $\bullet$
SNEFY 0.1350$\pm$0.0030 0.5981$\pm$0.0134 1.1283$\pm$0.0254 0.1209$\pm$0.0049 0.9191$\pm$0.0030 0.8098$\pm$0.0042
Table 3: The label distribution active learning performance on the Movie dataset.
Label Distribution Active Learning
In addition, we conduct experiments on label distribution active learning to evaluate SNEFY-LDL’s performance in uncertainty quantification. We choose the Movie dataset, randomly select 400 labeled training samples to form the initial labeled pool and take the remaining training samples as the unlabeled pool on each fold of cross-validation. The SNEFY-LDL model is first trained on the initial labeled pool. We then use different active learning strategies to pick up 200 informative samples from the unlabeled pool and query their labels. After augmenting the initial labeled pool with the 200 queried samples, we re-train the SNEFY-LDL model and evaluate its label distribution prediction performance on the test sets. Six different active learning strategies are compared:
  • Random (Zhan et al., 2022) randomly selects 200 samples from the unlabeled pool.
  • Kmeans (Zhdanov, 2019) selects samples close to the cluster centroids generated by the Kmeans clustering (MacKay, 2003) in feature space.
  • CoreSet (Sener and Savarese, 2018) selects the $k$-center samples (Har-Peled, 2011) as representative unlabeled samples, which is a variant of Kmeans.
  • Graph Density (Ebert et al., 2012) selects highly connected samples in the constructed KNN graph (Preparata and Shamos, 2012).
  • Dirichlet models the distribution of label distributions using a Dirichlet distribution (Ng et al., 2011) centered at predicted label distributions and selects samples with the largest differential entropies (Cover, 1999).
  • SNEFY-LDL uses importance sampling (Kloek and Van Dijk, 1978) to estimate differential entropies of the conditional distributions modeled by SNEFY-LDL and picks up samples having the largest differential entropies.
Among the six active learning strategies, Kmeans, CoreSet and Graph Density are representativeness based methods, which rely on only sample features by characterizing samples’ geometric properties in feature space, while Dirichlet and SNEFY-LDL are the uncertainty based methods we have contrived, which leverage the predictions of the initially trained SNEFY-LDL model.
Table 3 shows the performance of different active learning strategies, where the best performer is highlighted by boldface and $\bullet$ indicates that the annotated performer is inferior to the best performer at 0.05 significance level. As is shown in Table 3, SNEFY-LDL consistently achieves the best performance in terms of all metrics. By accurately evaluating label distribution prediction uncertainties, SNEFY-LDL can pick up more informative unlabeled samples than the naive uncertainty quantification strategy, Dirichlet, as well as the representativeness based active learning strategies, Kmeans, CoreSet and Graph Density, which are even sometimes inferior to the Random strategy.
Base Learner Bagging Cheby $\downarrow$ Clark $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
SA-BFGS Average 0.1178$\pm$0.0020$\bullet$ 0.3743$\pm$0.0068$\bullet$ 0.7948$\pm$0.0163$\bullet$ 0.0641$\pm$0.0022$\bullet$ 0.9370$\pm$0.0020$\bullet$ 0.8575$\pm$0.0028$\bullet$
Weighted 0.1137$\pm$0.0023 0.3625$\pm$0.0062 0.7686$\pm$0.0149 0.0604$\pm$0.0021 0.9406$\pm$0.0020 0.8624$\pm$0.0026
DF-LDL Average 0.1203$\pm$0.0017$\bullet$ 0.3762$\pm$0.0059$\bullet$ 0.8040$\pm$0.0146$\bullet$ 0.0657$\pm$0.0019$\bullet$ 0.9353$\pm$0.0018$\bullet$ 0.8557$\pm$0.0025$\bullet$
Weighted 0.1152$\pm$0.0024 0.3617$\pm$0.0070 0.7715$\pm$0.0168 0.0609$\pm$0.0024 0.9399$\pm$0.0023 0.8617$\pm$0.0030
LDL-SCL Average 0.1256$\pm$0.0020$\bullet$ 0.3828$\pm$0.0047$\bullet$ 0.8260$\pm$0.0118$\bullet$ 0.0699$\pm$0.0021$\bullet$ 0.9315$\pm$0.0018$\bullet$ 0.8519$\pm$0.0020$\bullet$
Weighted 0.1246$\pm$0.0023 0.3772$\pm$0.0048 0.8147$\pm$0.0114 0.0684$\pm$0.0022 0.9330$\pm$0.0019 0.8540$\pm$0.0020
LDL-LRR Average 0.1269$\pm$0.0021$\bullet$ 0.3966$\pm$0.0052$\bullet$ 0.8478$\pm$0.0131$\bullet$ 0.0730$\pm$0.0022$\bullet$ 0.9285$\pm$0.0019$\bullet$ 0.8478$\pm$0.0023$\bullet$
Weighted 0.1250$\pm$0.0020 0.3916$\pm$0.0044 0.8373$\pm$0.0106 0.0710$\pm$0.0020 0.9305$\pm$0.0018 0.8498$\pm$0.0019
Table 4: The label distribution ensemble learning performance on the SBU_3DFE dataset.
Label Distribution Ensemble Learning
We also conduct experiments on ensemble learning to further verify SNEFY-LDL’s ability in uncertainty quantification, by using Bagging (Breiman, 1996) as an exemplary ensemble learning paradigm. We choose the SBU_3DFE dataset, randomly select 50 samples from the training split of each fold of cross-validation for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of ensembled LDL models on the test sets. Four competitive LDL models are employed as base learners: SA-BFGS, DF-LDL, LDL-SCL and LDL-LRR, and two strategies are adopted to ensemble base learner predictions: 1) Average: aggregate the 25 base learner predictions with the uniform weight $1/25$, and 2) Weighted: weight each base learner prediction in proportion to its SNEFY-LDL conditional likelihood in an instance-wise manner.
Table 4 compares the two different ensemble learning strategies, where the best strategy is highlighted by boldface and $\bullet$ indicates that the annotated strategy is inferior to the best strategy at 0.05 significance level. From Table 4, we can find that the Weighted strategy is significantly better than Average in terms of all metrics. This implies that SNEFY-LDL provides an effective mechanism to quantify the prediction reliability/uncertainties of base learners.
Parameter Sensitivity Study
We also conduct additional experiments to investigate SNEFY-LDL’s sensitivity with regard to its four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training. The experiments demonstrate that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges. Please check the Appendix for more details.
7 Conclusion
In this paper, we propose a novel LDL paradigm: estimate the distribution of label distribution vectors in the probability simplex, which brings a bird’s-eye view on the relative significance of all possible label distributions. Through uncovering the underlying relationship between SNEFY and LDL, we develop the SNEFY-LDL model that can provide a tractable formulation of the conditional distribution of label distribution vectors, enjoying great expressivity and high computational efficiency. SNEFY-LDL admits closed-form expressions for the distribution’s mean, variance and covariance that can be computed quickly, making SNEFY-LDL able to provide real-time responses to the downstream tasks and real-world applications. Experiments on label distribution prediction, as well as label distribution active learning and ensemble learning demonstrate the great utility of SNEFY-LDL for label distribution modeling.
Appendix A Theorem Proofs
Theorem 2.
Let $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}}):\Delta^{L-1}\rightarrow\mathbb{R}^{L}$ by setting $D_{1}=L$, the activation function $\sigma$ be the exponential function $\exp$, the base measure $\mu_{1}(d\bm{\ell})=d\bm{\ell}$ be the Lebesgue measure. Under the condition that $\bm{W}_{1}>-1/2$ elementwise, the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ admits a closed form:
$\bm{k}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}},$ (10)
where $w_{1il}$ is the $il$-th element of matrix $\bm{W}_{1}$ and $\Gamma(\cdot)$ is the gamma function.
Proof.
According to Eq. (9),
$\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\int_{\Delta^{L-1}}\tilde{\bm{k}}_{\sigma,\bm{t}_{1},\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\ell},\bm{x})\mu_{1}(d\bm{\ell})$
$=$ $\int_{\Delta^{L-1}}\sigma(\bm{w}_{1i}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+b_{i})\cdot\sigma(\bm{w}_{1j}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{j})\mu_{1}(d\bm{\ell}).$
Given the setting $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$ and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}$ can be written as
[…]
$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}.$
As $\bm{W}_{1}>-1/2$ elementwise, $w_{1il}+w_{1jl}+1>0$. Assuming $\bm{\ell}$ follows a Dirichlet distribution (Ng, Tian, and Tang 2011) with parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{L})$, where $\alpha_{l}=w_{1il}+w_{1jl}+1>0$, its probability density, $\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})/{d\bm{\ell}}$, is in the form:
$\frac{\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})}{d\bm{\ell}}=\frac{1}{\mathrm{B}(\bm{\alpha})}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1},$
where $\mathrm{B}(\cdot)$ is the beta function. Considering the fact that $\int_{\Delta^{L-1}}\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})=1$,
$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1}d\bm{\ell}=\mathrm{B}(\alpha).$
That is to say
$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}=\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}.$
Therefore,
$\bm{k}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}.$
Theorem 3.
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (7) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
$\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{F}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$ (11)
where $\circ$ denotes Hadamard product, and $\bm{F}^{y_{r}}$ is a $n\times n$ matrix, whose $ij$th entry is
$F_{ij}^{y_{r}}=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})}.$ (12)
The conditional variance of $\bm{\ell}^{y_{r}}$, $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$, is
$\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{G}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}-\mathrm{E}^{2}[\bm{\ell}^{y_{r}}|\bm{x}],$ (13)
where $\bm{G}^{y_{r}}$ is a $n\times n$ matrix , with its $ij$th element being
$G^{y_{r}}_{ij}=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$ (14)
For two different labels $y_{r}$ and $y_{s}$, with $y_{r}\neq y_{s}$, the conditional covariance of $\bm{\ell}^{y_{r}}$ and $\bm{\ell}^{y_{s}}$, $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$, is
$\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{H}^{y_{r},y_{s}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]\cdot\mathrm{E}[\bm{\ell}^{y_{s}}|\bm{x}],$ (15)
where $\bm{H}^{y_{r},y_{s}}$ is a $n\times n$ matrix , with its $ij$th element being
$H^{y_{r},y_{s}}_{ij}=\frac{(1+w_{1ir}+w_{1jr})(1+w_{1is}+w_{1js})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$ (16)
Proof.
For any fixed function $\varphi(\bm{\ell})$ of $\bm{\ell}$, its expectation with regard to the conditional SNEFY distribution in Eq. (7), $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]$, can be computed as
$\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]=$ $\int_{\Delta^{L-1}}\varphi(\bm{\ell})\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$
$=$ $\int_{\Delta^{L-1}}\varphi(\bm{\ell})\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x}))}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}\mu_{1}(d\bm{\ell})$
$=$ $\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{\Phi}_{\bm{\Theta}}(\bm{x}))}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$
where $\bm{\Phi}_{\bm{\Theta}}(\bm{x})\in\mathbb{R}^{n\times n}$ is the elementwise integral:
$\bm{\Phi}_{\bm{\Theta}}(\bm{x})=\int_{\Delta^{L-1}}\varphi(\bm{\ell})\widetilde{\bm{K}_{\bm{\Theta}}}(\bm{\ell},\bm{x})\mu_{1}(d\bm{\ell}),$
whose $ij$th element is
$\bm{\phi}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=\int_{\Delta^{L-1}}\varphi(\bm{\ell})\tilde{\bm{k}}_{\sigma,\bm{t}_{1},\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{\ell},\bm{x})\mu_{1}(d\bm{\ell})$
$=\int_{\Delta^{L-1}}\varphi(\bm{\ell})\sigma(\bm{w}_{1i}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+b_{i})\cdot\sigma(\bm{w}_{1j}^{\top}\bm{t}_{1}(\bm{\ell})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{j})\mu_{1}(d\bm{\ell}).$
By setting $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, $\bm{\phi}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ can be written as
[…]
$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}$
As $\bm{W}_{1}>-1/2$ elementwise, $w_{1il}+w_{1jl}+1>0$. Assuming $\bm{\ell}$ follows a Dirichlet distribution with parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{L})$, where $\alpha_{l}=w_{1il}+w_{1jl}+1>0$, its probability density, $\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})/d\bm{\ell}$, is in the form:
$\frac{\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})}{d\bm{\ell}}=\frac{1}{\mathrm{B}(\bm{\alpha})}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1},$
where $\mathrm{B}(\cdot)$ is the beta function. Then, we have
$\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}$ $=\mathrm{B}(\bm{\alpha})\int_{{\Delta}^{L-1}}\varphi(\bm{\ell})\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})$
$=\mathrm{B}(\bm{\alpha})\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
$=\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}],$
where $\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$ is the expectation of $\varphi(\bm{\ell})$ with regard to the Dirichlet distribution parameterized by $\bm{w}_{1i}$ and $\bm{w}_{1j}$. Therefore,
$\bm{\phi}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})$
$\cdot$ $\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
$=$ $\bm{k}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}].$
By using $\bm{E}_{\varphi}$ to denote the $n\times n$ matrix whose $ij$th element is $\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$, we have
$\bm{\Phi}_{\bm{\Theta}}(\bm{x})=\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{E}_{\varphi},$
$\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{E}_{\varphi})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))}.$
For the Dirichlet distribution, $\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$ has closed forms for some moments. In particular, for $\varphi(\bm{\ell})=\bm{\ell}^{y_{r}}$, $\varphi(\bm{\ell})=(\bm{\ell}^{y_{r}})^{2}$, and $\varphi(\bm{\ell})=\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}}$ with $y_{r}\neq y_{s}$, we respectively have
$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})}.$
$\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$
$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(1+w_{1is}+w_{1js})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$
Finally, we have
$\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{F}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$
$\mathrm{E}[(\bm{\ell}^{y_{r}})^{2}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{G}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$
$\mathrm{E}[\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{H}^{y_{r},y_{s}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$
where $\bm{F}^{y_{r}}$, $\bm{G}^{y_{r}}$ and $\bm{H}^{y_{r},y_{s}}$ denote the $n\times n$ matrices whose $ij$th elements are $\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]$, $\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]$ and $\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}};\bm{w}_{1i},\bm{w}_{1j}]$ respectively.
$\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ and $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$ can be directly derived by using the identities that $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]=\mathrm{E}[(\bm{\ell}^{y_{r}})^{2}|\bm{x}]-\mathrm{E}^{2}[\bm{\ell}^{y_{r}}|\bm{x}]$ and $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]=\mathrm{E}[\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}}|\bm{x}]-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]\cdot\mathrm{E}[\bm{\ell}^{y_{s}}|\bm{x}]$. ∎
Appendix B Label Distribution Prediction Results
Tables 5-9 show the label distribution prediction performance measured by Chebyshev distance, Clark distance, Canberra metric, Kullback-Leibler divergence and Intersection, respectively, as the averaged scores on ten-fold cross-validation. On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline. We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level. The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline. The scores in Tables 5-9 consistently confirm that SNEFY-LDL is very competitive compared with the state-of-the-art baselines.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.1183 $\bullet$ 0.1193 $\bullet$ 0.1330 $\bullet$ 0.1163 $\spadesuit$ 0.1668 $\bullet$ 0.1135 $\circ$ 0.1444 $\bullet$ 0.1154 $\circ$ 0.1172 +0.0037
Natural_Scene 0.3104 $\spadesuit$ 0.3130 $\spadesuit$ 0.3563 $\bullet$ 0.3194 $\spadesuit$ 0.3023 $\spadesuit$ 0.3077 $\spadesuit$ 0.3603 $\bullet$ 0.3046 $\spadesuit$ 0.3135 +0.0112
SBU_3DFE 0.1300 $\bullet$ 0.1277 $\bullet$ 0.1394 $\bullet$ 0.1152 $\bullet$ 0.1197 $\bullet$ 0.1147 $\bullet$ 0.1303 $\bullet$ 0.1268 $\bullet$ 0.1110 -0.0037
SJAFFE 0.1158 $\bullet$ 0.0975 $\spadesuit$ 0.1202 $\bullet$ 0.0921 $\spadesuit$ 0.0862 $\circ$ 0.0810 $\circ$ 0.0868 $\circ$ 0.0831 $\circ$ 0.0997 +0.0188
Yeast_alpha 0.0140 $\bullet$ 0.0144 $\bullet$ 0.0140 $\bullet$ 0.0134 $\circ$ 0.0138 $\spadesuit$ 0.0134 $\circ$ 0.0135 $\circ$ 0.0134 $\circ$ 0.0137 +0.0003
Yeast_cdc 0.0168 $\spadesuit$ 0.0172 $\bullet$ 0.0169 $\bullet$ 0.0162 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0162 $\circ$ 0.0167 +0.0005
Yeast_cold 0.0529 $\bullet$ 0.0542 $\bullet$ 0.0538 $\bullet$ 0.0512 $\spadesuit$ 0.0535 $\bullet$ 0.0514 $\spadesuit$ 0.0543 $\bullet$ 0.0512 $\spadesuit$ 0.0518 +0.0006
Yeast_diau 0.0386 $\spadesuit$ 0.0384 $\spadesuit$ 0.0407 $\bullet$ 0.0369 $\circ$ 0.0376 $\spadesuit$ 0.0369 $\circ$ 0.0405 $\bullet$ 0.0370 $\circ$ 0.0382 +0.0013
Yeast_dtt 0.0381 $\bullet$ 0.0385 $\bullet$ 0.0370 $\spadesuit$ 0.0360 $\circ$ 0.0374 $\spadesuit$ 0.0365 $\spadesuit$ 0.0393 $\spadesuit$ 0.0360 $\circ$ 0.0369 +0.0009
Yeast_elu 0.0171 $\bullet$ 0.0173 $\bullet$ 0.0169 $\spadesuit$ 0.0163 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0163 $\circ$ 0.0168 +0.0005
Yeast_heat 0.0438 $\bullet$ 0.0440 $\bullet$ 0.0433 $\bullet$ 0.0422 $\spadesuit$ 0.0425 $\spadesuit$ 0.0420 $\circ$ 0.0435 $\spadesuit$ 0.0422 $\spadesuit$ 0.0425 +0.0006
Yeast_spo 0.0589 $\bullet$ 0.0629 $\bullet$ 0.0593 $\bullet$ 0.0584 $\spadesuit$ 0.0612 $\bullet$ 0.0603 $\bullet$ 0.0592 $\bullet$ 0.0583 $\spadesuit$ 0.0583 -0.0000
Yeast_spo5 0.0923 $\bullet$ 0.0952 $\bullet$ 0.0920 $\bullet$ 0.0913 $\spadesuit$ 0.0998 $\bullet$ 0.0973 $\bullet$ 0.1083 $\bullet$ 0.0912 $\spadesuit$ 0.0910 -0.0002
Yeast_spoem 0.0884 $\bullet$ 0.0907 $\bullet$ 0.0874 $\spadesuit$ 0.0868 $\spadesuit$ 0.0932 $\bullet$ 0.1043 $\bullet$ 0.2315 $\bullet$ 0.0865 $\spadesuit$ 0.0865 -0.0000
Table 5: The comparison of label distribution prediction performance measured by Chebyshev $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.5293 $\circ$ 0.5312 $\spadesuit$ 0.5636 $\bullet$ 0.5251 $\circ$ 0.7024 $\bullet$ 0.5107 $\circ$ 0.6230 $\bullet$ 0.5214 $\circ$ 0.5370 +0.0263
Natural_Scene 2.4441 $\circ$ 1.9084 $\circ$ 2.4878 $\bullet$ 2.4234 $\circ$ 2.4313 $\circ$ 2.4130 $\circ$ 2.3683 $\circ$ 2.4399 $\circ$ 2.4568 +0.5484
SBU_3DFE 0.4067 $\bullet$ 0.4015 $\bullet$ 0.4151 $\bullet$ 0.3710 $\bullet$ 0.3770 $\bullet$ 0.3624 $\bullet$ 0.4025 $\bullet$ 0.3975 $\bullet$ 0.3429 -0.0195
SJAFFE 0.4167 $\bullet$ 0.3477 $\spadesuit$ 0.4293 $\bullet$ 0.3589 $\spadesuit$ 0.3350 $\circ$ 0.3031 $\circ$ 0.3222 $\circ$ 0.3149 $\circ$ 0.3613 +0.0582
Yeast_alpha 0.2218 $\bullet$ 0.2261 $\bullet$ 0.2239 $\bullet$ 0.2099 $\circ$ 0.2166 $\spadesuit$ 0.2094 $\circ$ 0.2122 $\circ$ 0.2099 $\circ$ 0.2182 +0.0088
Yeast_cdc 0.2255 $\bullet$ 0.2317 $\bullet$ 0.2243 $\bullet$ 0.2158 $\circ$ 0.2173 $\circ$ 0.2157 $\circ$ 0.2179 $\circ$ 0.2159 $\circ$ 0.2219 +0.0062
Yeast_cold 0.1443 $\bullet$ 0.1477 $\bullet$ 0.1458 $\bullet$ 0.1396 $\spadesuit$ 0.1456 $\bullet$ 0.1405 $\spadesuit$ 0.1479 $\bullet$ 0.1396 $\spadesuit$ 0.1411 +0.0015
Yeast_diau 0.2096 $\bullet$ 0.2080 $\spadesuit$ 0.2178 $\bullet$ 0.2007 $\circ$ 0.2032 $\spadesuit$ 0.2001 $\circ$ 0.2213 $\bullet$ 0.2007 $\circ$ 0.2070 +0.0068
Yeast_dtt 0.1041 $\bullet$ 0.1042 $\bullet$ 0.1008 $\spadesuit$ 0.0982 $\circ$ 0.1017 $\spadesuit$ 0.0994 $\spadesuit$ 0.1065 $\spadesuit$ 0.0982 $\circ$ 0.1006 +0.0025
Yeast_elu 0.2114 $\bullet$ 0.2131 $\bullet$ 0.2073 $\spadesuit$ 0.1990 $\circ$ 0.2003 $\circ$ 0.1985 $\circ$ 0.2014 $\circ$ 0.1990 $\circ$ 0.2077 +0.0092
Yeast_heat 0.1893 $\bullet$ 0.1908 $\bullet$ 0.1876 $\bullet$ 0.1826 $\spadesuit$ 0.1844 $\spadesuit$ 0.1815 $\circ$ 0.1882 $\bullet$ 0.1826 $\spadesuit$ 0.1839 +0.0024
Yeast_spo 0.2530 $\bullet$ 0.2665 $\bullet$ 0.2533 $\bullet$ 0.2498 $\spadesuit$ 0.2594 $\bullet$ 0.2552 $\bullet$ 0.2530 $\bullet$ 0.2497 $\spadesuit$ 0.2494 -0.0003
Yeast_spo5 0.1861 $\bullet$ 0.1915 $\bullet$ 0.1857 $\bullet$ 0.1840 $\spadesuit$ 0.1999 $\bullet$ 0.1951 $\bullet$ 0.2145 $\bullet$ 0.1838 $\spadesuit$ 0.1833 -0.0005
Yeast_spoem 0.1313 $\bullet$ 0.1350 $\bullet$ 0.1299 $\spadesuit$ 0.1292 $\spadesuit$ 0.1392 $\bullet$ 0.1541 $\bullet$ 0.3628 $\bullet$ 0.1287 $\spadesuit$ 0.1288 +0.0001
Table 6: The comparison of label distribution prediction performance measured by Clark distance $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 1.0116 $\spadesuit$ 1.0205 $\spadesuit$ 1.0588 $\bullet$ 1.0055 $\spadesuit$ 1.4036 $\bullet$ 0.9796 $\circ$ 1.1795 $\bullet$ 0.9970 $\circ$ 1.0154 +0.0357
Natural_Scene 6.7001 $\spadesuit$ 4.5465 $\circ$ 6.9670 $\bullet$ 6.6520 $\circ$ 6.6519 $\spadesuit$ 6.5907 $\circ$ 6.5539 $\circ$ 6.6785 $\circ$ 6.7202 +2.1737
SBU_3DFE 0.8713 $\bullet$ 0.8306 $\bullet$ 0.9055 $\bullet$ 0.7827 $\bullet$ 0.8035 $\bullet$ 0.7638 $\bullet$ 0.8580 $\bullet$ 0.8485 $\bullet$ 0.7347 -0.0291
SJAFFE 0.8604 $\bullet$ 0.7148 $\spadesuit$ 0.9001 $\bullet$ 0.7326 $\spadesuit$ 0.6820 $\circ$ 0.6188 $\circ$ 0.6648 $\circ$ 0.6421 $\circ$ 0.7473 +0.1285
Yeast_alpha 0.7237 $\bullet$ 0.7388 $\bullet$ 0.7327 $\bullet$ 0.6818 $\circ$ 0.7051 $\spadesuit$ 0.6802 $\circ$ 0.6911 $\circ$ 0.6819 $\circ$ 0.7128 +0.0325
Yeast_cdc 0.6794 $\bullet$ 0.7001 $\bullet$ 0.6733 $\spadesuit$ 0.6473 $\circ$ 0.6524 $\circ$ 0.6472 $\circ$ 0.6540 $\circ$ 0.6475 $\circ$ 0.6673 +0.0202
Yeast_cold 0.2487 $\bullet$ 0.2550 $\bullet$ 0.2515 $\bullet$ 0.2403 $\spadesuit$ 0.2513 $\bullet$ 0.2424 $\spadesuit$ 0.2558 $\bullet$ 0.2403 $\spadesuit$ 0.2430 +0.0027
Yeast_diau 0.4501 $\bullet$ 0.4476 $\spadesuit$ 0.4671 $\bullet$ 0.4308 $\circ$ 0.4367 $\spadesuit$ 0.4300 $\circ$ 0.4804 $\bullet$ 0.4308 $\circ$ 0.4449 +0.0149
Yeast_dtt 0.1796 $\bullet$ 0.1789 $\bullet$ 0.1732 $\spadesuit$ 0.1689 $\circ$ 0.1750 $\spadesuit$ 0.1711 $\spadesuit$ 0.1832 $\spadesuit$ 0.1689 $\circ$ 0.1731 +0.0043
Yeast_elu 0.6236 $\bullet$ 0.6281 $\bullet$ 0.6084 $\spadesuit$ 0.5831 $\circ$ 0.5876 $\circ$ 0.5820 $\circ$ 0.5905 $\circ$ 0.5832 $\circ$ 0.6116 +0.0296
Yeast_heat 0.3780 $\bullet$ 0.3834 $\bullet$ 0.3739 $\bullet$ 0.3641 $\circ$ 0.3688 $\spadesuit$ 0.3626 $\circ$ 0.3763 $\bullet$ 0.3642 $\circ$ 0.3670 +0.0044
Yeast_spo 0.5199 $\bullet$ 0.5489 $\bullet$ 0.5223 $\bullet$ 0.5131 $\spadesuit$ 0.5346 $\bullet$ 0.5245 $\bullet$ 0.5218 $\bullet$ 0.5129 $\spadesuit$ 0.5143 +0.0014
Yeast_spo5 0.2856 $\bullet$ 0.2945 $\bullet$ 0.2847 $\bullet$ 0.2826 $\spadesuit$ 0.3082 $\bullet$ 0.3003 $\bullet$ 0.3344 $\bullet$ 0.2824 $\spadesuit$ 0.2820 -0.0004
Yeast_spoem 0.1828 $\bullet$ 0.1878 $\bullet$ 0.1807 $\spadesuit$ 0.1797 $\spadesuit$ 0.1934 $\bullet$ 0.2149 $\bullet$ 0.4967 $\bullet$ 0.1791 $\spadesuit$ 0.1792 +0.0001
Table 7: The comparison of label distribution prediction performance measured by Canberra metric $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.1011 $\bullet$ 0.1093 $\bullet$ 0.1113 $\bullet$ 0.1002 $\bullet$ 0.2357 $\bullet$ 0.0975 $\spadesuit$ 0.1380 $\bullet$ 0.0981 $\spadesuit$ 0.0984 +0.0009
Natural_Scene 0.7608 $\spadesuit$ 1.0936 $\bullet$ 1.0466 $\bullet$ 0.8277 $\spadesuit$ 0.7832 $\spadesuit$ 0.7608 $\spadesuit$ 1.0478 $\bullet$ 0.7445 $\spadesuit$ 0.7704 +0.0259
SBU_3DFE 0.0774 $\bullet$ 0.0803 $\bullet$ 0.0860 $\bullet$ 0.0629 $\bullet$ 0.0651 $\bullet$ 0.0620 $\bullet$ 0.0776 $\bullet$ 0.0730 $\bullet$ 0.0578 -0.0042
SJAFFE 0.0701 $\bullet$ 0.0534 $\spadesuit$ 0.0740 $\bullet$ 0.0505 $\spadesuit$ 0.0446 $\circ$ 0.0375 $\circ$ 0.0419 $\circ$ 0.0391 $\circ$ 0.0523 +0.0148
Yeast_alpha 0.0060 $\bullet$ 0.0063 $\bullet$ 0.0061 $\bullet$ 0.0055 $\circ$ 0.0058 $\spadesuit$ 0.0055 $\circ$ 0.0056 $\circ$ 0.0055 $\circ$ 0.0059 +0.0004
Yeast_cdc 0.0075 $\bullet$ 0.0079 $\bullet$ 0.0076 $\bullet$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0073 +0.0004
Yeast_cold 0.0128 $\bullet$ 0.0136 $\bullet$ 0.0132 $\bullet$ 0.0122 $\circ$ 0.0132 $\bullet$ 0.0123 $\spadesuit$ 0.0137 $\bullet$ 0.0122 $\circ$ 0.0124 +0.0002
Yeast_diau 0.0142 $\bullet$ 0.0145 $\spadesuit$ 0.0151 $\bullet$ 0.0131 $\circ$ 0.0138 $\spadesuit$ 0.0132 $\circ$ 0.0165 $\bullet$ 0.0131 $\circ$ 0.0138 +0.0007
Yeast_dtt 0.0069 $\bullet$ 0.0070 $\bullet$ 0.0065 $\spadesuit$ 0.0063 $\circ$ 0.0068 $\spadesuit$ 0.0064 $\spadesuit$ 0.0072 $\spadesuit$ 0.0063 $\circ$ 0.0065 +0.0002
Yeast_elu 0.0069 $\bullet$ 0.0071 $\bullet$ 0.0066 $\spadesuit$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0066 +0.0005
Yeast_heat 0.0135 $\bullet$ 0.0139 $\bullet$ 0.0133 $\bullet$ 0.0126 $\circ$ 0.0130 $\spadesuit$ 0.0125 $\circ$ 0.0133 $\spadesuit$ 0.0126 $\circ$ 0.0129 +0.0004
Yeast_spo 0.0251 $\spadesuit$ 0.0289 $\bullet$ 0.0258 $\bullet$ 0.0246 $\circ$ 0.0271 $\bullet$ 0.0259 $\bullet$ 0.0257 $\bullet$ 0.0246 $\circ$ 0.0251 +0.0005
Yeast_spo5 0.0298 $\bullet$ 0.0326 $\bullet$ 0.0297 $\bullet$ 0.0292 $\spadesuit$ 0.0351 $\bullet$ 0.0338 $\bullet$ 0.0409 $\bullet$ 0.0292 $\spadesuit$ 0.0290 -0.0002
Yeast_spoem 0.0253 $\bullet$ 0.0274 $\bullet$ 0.0247 $\spadesuit$ 0.0245 $\spadesuit$ 0.0301 $\bullet$ 0.0346 $\bullet$ 0.1475 $\bullet$ 0.0245 $\spadesuit$ 0.0243 -0.0002
Table 8: The comparison of label distribution prediction performance measured by Kullback-Leibler divergence $\downarrow$.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.8324 $\spadesuit$ 0.8294 $\bullet$ 0.8196 $\bullet$ 0.8336 $\spadesuit$ 0.7572 $\bullet$ 0.8380 $\circ$ 0.7953 $\bullet$ 0.8353 $\circ$ 0.8333 -0.0048
Natural_Scene 0.5549 $\spadesuit$ 0.5629 $\circ$ 0.4153 $\bullet$ 0.5491 $\spadesuit$ 0.5536 $\spadesuit$ 0.5566 $\spadesuit$ 0.5294 $\spadesuit$ 0.5595 $\circ$ 0.5343 -0.0286
SBU_3DFE 0.8436 $\bullet$ 0.8484 $\bullet$ 0.8380 $\bullet$ 0.8600 $\bullet$ 0.8559 $\bullet$ 0.8624 $\bullet$ 0.8456 $\bullet$ 0.8477 $\bullet$ 0.8685 +0.0061
SJAFFE 0.8523 $\bullet$ 0.8768 $\spadesuit$ 0.8466 $\bullet$ 0.8786 $\spadesuit$ 0.8865 $\circ$ 0.8959 $\circ$ 0.8876 $\circ$ 0.8928 $\circ$ 0.8729 -0.0229
Yeast_alpha 0.9600 $\bullet$ 0.9592 $\bullet$ 0.9595 $\bullet$ 0.9624 $\circ$ 0.9611 $\spadesuit$ 0.9624 $\circ$ 0.9618 $\circ$ 0.9624 $\circ$ 0.9606 -0.0018
Yeast_cdc 0.9552 $\bullet$ 0.9539 $\bullet$ 0.9557 $\spadesuit$ 0.9574 $\circ$ 0.9571 $\circ$ 0.9574 $\circ$ 0.9570 $\circ$ 0.9574 $\circ$ 0.9561 -0.0014
Yeast_cold 0.9387 $\bullet$ 0.9371 $\bullet$ 0.9379 $\bullet$ 0.9408 $\spadesuit$ 0.9380 $\bullet$ 0.9403 $\spadesuit$ 0.9368 $\bullet$ 0.9408 $\spadesuit$ 0.9401 -0.0007
Yeast_diau 0.9375 $\bullet$ 0.9379 $\spadesuit$ 0.9350 $\bullet$ 0.9403 $\circ$ 0.9394 $\spadesuit$ 0.9404 $\circ$ 0.9334 $\bullet$ 0.9403 $\circ$ 0.9383 -0.0021
Yeast_dtt 0.9556 $\bullet$ 0.9559 $\bullet$ 0.9572 $\spadesuit$ 0.9583 $\circ$ 0.9568 $\spadesuit$ 0.9578 $\spadesuit$ 0.9547 $\spadesuit$ 0.9583 $\circ$ 0.9573 -0.0011
Yeast_elu 0.9560 $\bullet$ 0.9557 $\bullet$ 0.9571 $\spadesuit$ 0.9589 $\circ$ 0.9585 $\circ$ 0.9589 $\circ$ 0.9583 $\circ$ 0.9589 $\circ$ 0.9568 -0.0021
Yeast_heat 0.9379 $\bullet$ 0.9371 $\bullet$ 0.9386 $\bullet$ 0.9402 $\circ$ 0.9395 $\spadesuit$ 0.9405 $\circ$ 0.9382 $\bullet$ 0.9402 $\circ$ 0.9397 -0.0008
Yeast_spo 0.9144 $\bullet$ 0.9094 $\bullet$ 0.9140 $\bullet$ 0.9155 $\spadesuit$ 0.9118 $\bullet$ 0.9135 $\bullet$ 0.9140 $\bullet$ 0.9156 $\spadesuit$ 0.9153 -0.0002
Yeast_spo5 0.9077 $\bullet$ 0.9048 $\bullet$ 0.9080 $\bullet$ 0.9087 $\spadesuit$ 0.9002 $\bullet$ 0.9027 $\bullet$ 0.8917 $\bullet$ 0.9088 $\spadesuit$ 0.9090 +0.0002
Yeast_spoem 0.9116 $\bullet$ 0.9093 $\bullet$ 0.9126 $\spadesuit$ 0.9132 $\spadesuit$ 0.9068 $\bullet$ 0.8957 $\bullet$ 0.7685 $\bullet$ 0.9135 $\spadesuit$ 0.9135 +0.0000
Table 9: The comparison of label distribution prediction performance measured by Intersection $\uparrow$.
Appendix C Parameter Sensitivity Study
On the Movie and SBU_3DFE datasets, we take turns to study the sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training, by varying the studied hyperparameter in a predefined range and fixing the remaining three as default values at each turn. Figure 1 plots the performance change of SNEFY-LDL measured by Cosine coefficient with the varying values of the four hyperparameters. From Figure 1, we can find that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges.
(a) $n$
(b) $m$
(c) batch size
(d) epoch number
Figure 1: Parameter sensitivity.

### v2

@@
 Label distribution learning (LDL) provides a framework wherein a distribution over categories rather than a single category is predicted, with the aim of addressing ambiguity in labeled data.
-Existing research on LDL mainly focuses on the task of point estimation, i.e., pinpointing an optimal distribution in the probability simplex conditioned on the input sample.
-In this paper, we estimate a probability distribution of all possible label distributions over the simplex, by unleashing the expressive power of the recently introduced Squared Neural Family (SNEFY).
-With the modeled distribution, label distribution prediction can be achieved by performing the expectation operation to estimate the mean of the distribution of label distributions.
-Moreover, more information about the label distribution can be inferred, such as the prediction reliability and uncertainties.
-We conduct extensive experiments on the label distribution prediction task, showing that our distribution modeling based method can achieve very competitive label distribution prediction performance compared with the state-of-the-art baselines.
-Additional experiments on active learning and ensemble learning demonstrate that our probabilistic approach can effectively boost the performance in these settings, by accurately estimating the prediction reliability and uncertainties.
+Existing research on LDL mainly focuses on the task of point estimation, i.e., finding an optimal distribution in the probability simplex conditioned on the given sample.
+In this paper, we propose a novel label distribution learning model SNEFY-LDL, which estimates a probability distribution of all possible label distributions over the simplex, by unleashing the expressive power of the recently introduced Squared Neural Family (SNEFY), a new class of tractable probability models.
+As a way to summarize the fitted model, we derive the closed-form label distribution mean, variance and covariance conditioned on the given sample, which can be used to predict the ground-truth label distributions, construct label distribution confidence intervals, and measure the correlations between different labels.
+Moreover, more information about the label distribution prediction uncertainties can be acquired from the modeled probability density function.
+Extensive experiments on conformal prediction, active learning and ensemble learning are conducted, verifying SNEFY-LDL’s great effectiveness in LDL uncertainty quantification.
+The source code of this paper is available at https://github.com/daokunzhang/SNEFY-LDL.
 1 Introduction
-Label distribution learning (LDL) is a technique which handles ambiguity in multi-class classification, by utilising simplex-valued rather than categorical-valued labels in training data.
-Unlike traditional multi-class and multi-label learning paradigms, which assign a deterministic label prediction to instances, LDL corresponds to the question “How well does each of the labels describe an instance?”, by using a discrete probability distribution to characterize the instance’s affinity to all candidate categories.
+Label distribution learning (LDL) is a technique which handles ambiguity in multi-class classification, by utilizing simplex-valued rather than categorical-valued labels in training data.
+Unlike traditional multi-class and multi-label learning paradigms, which assign a deterministic label prediction to instances, LDL corresponds to the question "How well does each of the labels describe an instance?", by using a discrete probability distribution to characterize each label’s composition ratio in jointly describing the given instance.
 For example, when we predict the functionality of a district in a city, we might predict a result such as: the district has 20% functionality for business, 40% functionality for entertainment, and 40% functionality for education.
-Many LDL algorithms have been proposed to directly predict label distribution vectors from instance features, by adapting machine learning algorithms designed for “hard” label prediction to the “soft” label prediction setting.
+Many LDL algorithms have been proposed to directly predict label distribution vectors from instance features, by adapting machine learning algorithms designed for "hard" label prediction to the "soft" label prediction setting.
 Though a discrete distribution among candidate labels is predicted, existing LDL algorithms still operate at the level of point estimation, i.e., they search for a single point on a probability simplex (the set of all possible label distributions) for each given instance.
@@
 An additional bonus of the distribution modeling is the ability to quantify the prediction reliability and uncertainties, which not only facilitates reliable model deployment in real-world safety critical applications, but is also essential to various reliability/uncertainty-aware tasks, like pseudo labeling, active learning, and ensemble learning.
-Contributions
-In this paper, we propose a novel LDL framework, SNEFY-LDL, by unleashing the probability modeling power of the recently proposed Squared Neural Family (SNEFY) (Tsuchida et al., 2023), a new class of tractable probability models.
-By restricting the support set of SNEFY to a probability simplex, SNEFY-LDL provides an expressive conditional distribution modeling of the label distribution vector given the input sample.
+Contributions.
+In this paper, we propose a novel LDL framework, SNEFY-LDL, by unleashing the probability modeling power of the recently introduced Squared Neural Family (SNEFY) [Tsuchida et al., 2023], a new class of tractable probability models.
+By restricting the support set of SNEFY to a probability simplex, SNEFY-LDL constructs an expressive multimodal distribution modeling of the label distribution vector conditioned the given sample.
 The conditional distribution model has a closed-form normalizing constant, guaranteeing computational tractability.
 In this way, model parameters can be learned efficiently by maximizing the conditional likelihoods of training samples with stochastic gradient descent.
-Through computing the expectation of the conditional distribution, we can obtain the closed-form distribution mean, as a way to summarize the fitted model and predict label distributions for unlabeled samples.
-However, this is only one way to use the fitted model and the probability density values can be used to evaluate the reliability of label distribution predictions.
-We conduct extensive experiments on fourteen real-world datasets to evaluate the label distribution prediction performance of the proposed SNEFY-LDL model.
-Experimental results show that SNEFY-LDL can achieve very competitive label distribution prediction performance, compared with the state-of-the-art LDL baselines.
-In addition, label distribution active learning and ensemble learning experiments are carried out to verify the efficacy of SNEFY-LDL in quantifying prediction uncertainties.
+As a way to summarize the fitted model, we derive the closed-form label distribution mean, variance and covariance conditioned on the given sample, which can be used to predict the ground-truth label distribution, construct label distribution confidence intervals, and measure the correlations between different labels.
+However, the fitted model is not limited to these usecases, and the probability density values can be used to directly evaluate the reliability of label distribution predictions.
+We conduct extensive experiments on label distribution conformal prediction, active learning and ensemble learning to verify the efficacy of SNEFY-LDL in quantifying prediction uncertainties.
+For the task of conformal prediction, we use the SNEFY-LDL’s closed-form conditional mean and variance of label distribution predictions to construct a confidence interval for each label’s composition ratio in describing the given instance and calibrate the confidence interval through conformal prediction [Angelopoulos and Bates, 2021].
+Experimental results show that the confidence intervals constructed by the SNEFY-LDL model have greater adaptivity than the confidence intervals constructed by the naive Dirichlet distribution.
 The max-entropy principle is used to achieve active learning with the estimated SNEFY-LDL entropy, i.e., select the most informative unlabeled samples with the largest entropy values, query their labels and augment training samples, to attain the largest performance gain of the re-trained LDL model.
@@
 2 Related Work
-LDL is first proposed by (Geng et al., 2013) to solve the facial age estimation problem with insufficient training samples.
-Since then, a series of LDL algorithms have been developed, which can be divided into three categories: Problem Transformation (PT), Algorithm Adaptation (AA) and Specialized Algorithms (SA) (Geng, 2016).
-Problem Transformation
-PT (Geng, 2016) transforms the LDL problem into the single-label classification problem, by decomposing each training sample assigned with a label distribution into a set of duplicate training samples.
+LDL is first proposed by Geng et al.
+[2013] to solve the facial age estimation problem.
+Since then, a series of LDL algorithms have been developed, which are mainly in three categories: Problem Transformation (PT), Algorithm Adaptation (AA) and Specialized Algorithms (SA) [Geng, 2016].
+Problem Transformation.
+PT [Geng, 2016] transforms the LDL problem into the single-label classification problem, by decomposing each training sample assigned with a label distribution into a set of duplicate training samples.
 Each of them is assigned with a different label and accounts for a partial sample in proportion to the label probability value, and is then used to train the single-label classifiers.
 The label likelihoods predicted by the single-label classifiers are then aggregated to form the final prediction of label distributions.
-PT-Bayes (Geng, 2016) and PT-SVM (Geng, 2016; Geng and Hou, 2015) transform the LDL problem into the single-label multi-class classification problem and respectively employ Bayes and SVM as the single-label classifiers.
-DF-LDL (González et al., 2021a) decomposes the label distribution prediction task into a number of one-versus-one binary classification tasks, and fuses the binary classification likelihoods to form the final label distribution prediction.
-Algorithm Adaptation
-AA (Geng, 2016) adapts traditional single-label classification models into the LDL setting, by leveraging the models’ compatibility in outputting a soft label distribution vector.
-Derived from the K Nearest Neighbor (KNN) algorithm (Wu et al., 2008), AA-KNN (Geng, 2016) predicts samples’ label distributions by averaging the label distributions of their $k$ nearest neighbors in feature space.
-AA-BP (Geng, 2016) constructs a three-layer neural network and adopts the softmax function as the activation of the output layer, making the neural network naturally produce a label distribution for each example.
-The neural network is trained by minimizing the sum of squared errors between the model output label distributions and the ground-truth label distributions.
-Specialized Algorithms
-SA (Geng, 2016) designs algorithms from scratch to directly solve the LDL problem.
-SA-IIS (Geng et al., 2013) and SA-BFGS (Geng, 2016) use the maximum entropy model to parameterize label distributions.
-They are trained by minimizing the Kullback-Leibler (KL) divergence between the model output and ground-truth label distributions, where Improved Iterative Scaling (IIS) (Malouf, 2002) and BFGS (Nocedal and Wright, 1999) are respectively leveraged by SA-IIS and SA-BFGS as optimizers.
-CPNN (Geng et al., 2013) uses a neural network to parameterize the joint probability distribution between sample features and labels following Modha’s probability distribution formulation (Modha and Fainman, 1994).
-BCPNN (Yang et al., 2017) and ACPNN (Yang et al., 2017) then improve on CPNN through leveraging binary label encoding and augmenting training samples respectively.
-LDLF (Shen et al., 2017) employs differentiable decision trees (Kontschieder et al., 2015) to model label distributions and KL divergence is used to design the learning objective.
-LDL-SCL (Jia et al., 2021) forces the label distributions of samples located closely in feature space to be similar to each other.
-LDL-LRR (Jia et al., 2023b) and LDL-DPA (Jia et al., 2023a) maintain the relative importance ranking between different labels in label distribution modeling, by penalizing a label importance ranking loss in their learning objectives.
-Extensions
-In addition, LDL has been extended to other tasks, like label enhancement (recovering label distributions from the one-hot single-label assignments) (Xu et al., 2019b, 2020; Zheng et al., 2023), multi-class classification (Wang and Geng, 2019, 2021b, 2021a), learning with incomplete supervision (Xu and Zhou, 2017), oversampling (González et al., 2021b), ordinal LDL (Wen et al., 2023), semi-supervised learning (Xie et al., 2023), and label calibration (He et al., 2024).
-Furthermore, LDL has been applied to solve numerous real-world problems, including facial age estimation (Geng et al., 2013), facial emotion recognition (Chen et al., 2020), head pose estimation (Xu et al., 2019a), crowd opinion prediction (Geng and Hou, 2015), emphasis selection (Shirani et al., 2019), lesion counting (Wu et al., 2019), and urban function distribution (Huang et al., 2023).
-It is worth noting that there is a related topic in the classical statistics literature, termed compositional data analysis (Greenacre, 2021), with a broad range of applications including geochemistry (e.g.
+PT-Bayes [Geng, 2016] and PT-SVM [Geng and Hou, 2015] transform the LDL problem into the single-label multi-class classification problem and respectively employ Bayes and SVM as the single-label classifiers.
+DF-LDL [González et al., 2021a] decomposes the label distribution prediction task into a number of one-versus-one binary classification tasks, and fuses the binary classification likelihoods to form the final label distribution predictions.
+Algorithm Adaptation.
+AA [Geng, 2016] adapts traditional single-label classification models into the LDL setting, by leveraging the models’ compatibility in outputting a soft label distribution vector.
+Derived from the K Nearest Neighbor (KNN) algorithm [Wu et al., 2008], AA-KNN [Geng, 2016] predicts samples’ label distributions by averaging the label distributions of their $k$ nearest neighbors in feature space.
+AA-BP [Geng, 2016] constructs a three-layer neural network and adopts the softmax function as the activation of the output layer, making the neural network naturally produce a label distribution for each example.
+The neural network is trained by minimizing the sum of squared errors between the model-output label distributions and the ground-truth label distributions.
+Specialized Algorithms.
+SA [Geng, 2016] designs algorithms from scratch to directly solve the LDL problem.
+SA-IIS [Geng et al., 2013] and SA-BFGS [Geng, 2016] use the maximum entropy model to parameterize label distributions.
+They are trained by minimizing the Kullback-Leibler (KL) divergence between the model-output and ground-truth label distributions, where Improved Iterative Scaling (IIS) [Malouf, 2002] and BFGS [Nocedal and Wright, 1999] are respectively leveraged by SA-IIS and SA-BFGS as optimizers.
+CPNN [Geng et al., 2013] uses a neural network to parameterize the joint probability distribution between sample features and labels following Modha’s probability distribution formulation [Modha and Fainman, 1994].
+BCPNN [Yang et al., 2017] and ACPNN [Yang et al., 2017] then improve on CPNN through leveraging binary label encoding and augmenting training samples respectively.
+LDLF [Shen et al., 2017] employs differentiable decision trees [Kontschieder et al., 2015] to model label distributions and KL divergence is used to design the learning objective.
+LDL-SCL [Jia et al., 2021] forces the label distributions of samples located closely in feature space to be similar to each other.
+LDL-LRR [Jia et al., 2023b] and LDL-DPA [Jia et al., 2023a] maintain the relative importance ranking between different labels in label distribution modeling, by penalizing a label importance ranking loss in their learning objectives.
+Extensions.
+In addition, LDL has been extended to other tasks, like label enhancement  [Xu et al., 2019b, 2020, Zheng et al., 2023], multi-class classification [Wang and Geng, 2019, 2021b, 2021a], learning with incomplete supervision [Xu and Zhou, 2017], oversampling [González et al., 2021b], ordinal LDL [Wen et al., 2023], semi-supervised learning [Xie et al., 2023], and label calibration [He et al., 2024].
+Furthermore, LDL has been applied to solve numerous real-world problems, including facial age estimation [Geng et al., 2013], facial emotion recognition [Chen et al., 2020], head pose estimation [Xu et al., 2019a], crowd opinion prediction [Geng and Hou, 2015], emphasis selection [Shirani et al., 2019], lesion counting [Wu et al., 2019], and urban functionality prediction [Huang et al., 2023].
+It is worth noting that there is a related topic in the classical statistics literature, termed compositional data analysis [Greenacre, 2021], with a broad range of applications including geochemistry (e.g.
 labels correspond to mineral compositions) and ecology (e.g.
@@
 In addition, each sample $\bm{x}\in\mathcal{X}$ is described by a $L$-dimensional label distribution vector $\bm{\ell}_{\bm{x}}\in\mathbb{R}^{L}$ that takes values in the $(L-1)$-simplex $\Delta^{L-1}$, corresponding to a set of $L$ given labels $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$.
-The $l$-th entry $\bm{\ell}_{\bm{x}}^{y_{l}}$ of the label distribution vector $\bm{\ell}_{\bm{x}}$ corresponds to the probability of sample $\bm{x}$ belonging to the $l$-th class $y_{l}$, i.e., $\bm{\ell}_{\bm{x}}^{y_{l}}=\mathrm{P}(y_{l}|\bm{x})\geq 0$, satisfying the constraint that $\sum_{l=1}^{L}\bm{\ell}_{\bm{x}}^{y_{l}}=1$.
+The $l$-th entry $\bm{\ell}_{\bm{x}}^{y_{l}}$ of the label distribution vector $\bm{\ell}_{\bm{x}}$ corresponds to the composition ratio of the $l$-th label $y_{l}$ in describing $\bm{x}$, satisfying the constraint that $\sum_{l=1}^{L}\bm{\ell}_{\bm{x}}^{y_{l}}=1$.
 With the label distribution observations of training samples, our objective is to model the probability distribution of the label distribution vector $\bm{\ell}\in\Delta^{L-1}$ conditioned on any input sample $\bm{x}\in\mathbb{R}^{d}$, $\mathrm{P}(d\bm{\ell}|\bm{x})$.
-For any sample $\bm{x}\in\mathbb{R}^{d}$ with an unknown label distribution, one can associate to it a label distribution vector $\bm{\ell}^{*}_{\bm{x}}$ obtained as the mean of the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x})$:
-$\bm{\ell}^{*}_{\bm{x}}=\int_{\Delta^{L-1}}\bm{\ell}\,\mathrm{P}(d\bm{\ell}|\bm{x}),$ (1)
-but the fitted model $\mathrm{P}(d\bm{\ell}|\bm{x})$ can serve many other purposes.
 4 Preliminaries on SNEFY
@@
 where $\bm{t}(\cdot):\mathcal{Z}\rightarrow\mathbb{R}^{D}$ is the sufficient statistic, $\sigma$ is the activation function, $\bm{W}\in\mathbb{R}^{n\times D}$ and $\bm{b}\in\mathbb{R}^{n}$ are respectively the weight matrix and bias vector at the hidden layer of neural network $\bm{f}$, and $\bm{V}\in\mathbb{R}^{m\times n}$ are $\bm{f}$’s readout parameters.
-$\bm{\Theta}=(\bm{W},\bm{b})\in\mathbb{R}^{n\times(D+1)}$ denotes the concatenation of $\bm{W}$ and $\bm{b}$ and its $i$th row is denoted as $\bm{\theta}_{i}=(\bm{w}_{i},b_{i})\in\mathbb{R}^{D+1}$.
+$\bm{\Theta}=(\bm{W},\bm{b})\in\mathbb{R}^{n\times(D+1)}$ is the concatenation of $\bm{W}$ and $\bm{b}$ and its $i$th row is denoted as $\bm{\theta}_{i}=(\bm{w}_{i},b_{i})\in\mathbb{R}^{D+1}$, where $\bm{w}_{i}\in\mathbb{R}^{D}$ is the $i$th row of $\bm{W}$ and $b_{i}$ is the $i$th element of $\bm{b}$.
 The distribution $\mathrm{P}(d\bm{z};\bm{V},\bm{\Theta})$ in Eq. (1) admits a more concise formulation
@@
 $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})=\int_{\mathcal{Z}}\tilde{\bm{k}}_{\sigma,\bm{t}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{z})\mu(d\bm{z}).$ (4)
-Under varying choices of the activation function $\sigma$, sufficient statistic $\bm{t}$, and the base measure $\mu$, $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})$ is able to be computed in closed form (see Table 1 of (Tsuchida et al., 2023)) in $\mathcal{O}(D)$.
+Under varying choices of the activation function $\sigma$, sufficient statistic $\bm{t}$, and the base measure $\mu$, $\bm{k}_{\sigma,\bm{t},\mu}(\bm{\theta}_{i},\bm{\theta}_{j})$ is able to be computed in closed form (see Table 1 of [Tsuchida et al., 2023]) in $\mathcal{O}(D)$.
 This makes SNEFY a tractable probability distribution model, with great expressivity and computational efficiency.
@@
 Theorem 1.
-(Tsuchida et al., 2023) Let $\mathtt{z}=(\mathtt{z}_{1},\mathtt{z}_{2})$ jointly follow a SNEFY distribution, with support set $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, sufficient statistic $\bm{t}$, activation function $\sigma$, base measure $\mu$, as well as parameters $\bm{V}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})$.
+[Tsuchida et al., 2023] Let $\mathtt{z}=(\mathtt{z}_{1},\mathtt{z}_{2})$ jointly follow a SNEFY distribution, with support set $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}$, sufficient statistic $\bm{t}$, activation function $\sigma$, base measure $\mu$, as well as parameters $\bm{V}$ and $\bm{\Theta}=([\bm{W}_{1},\bm{W}_{2}],\bm{b})$.
 Assume that $\mu(d\bm{z})=\mu_{1}(d\bm{z}_{1})\mu_{2}(d\bm{z}_{2})$ and $\bm{t}(\bm{z})=(\bm{t}_{1}(\bm{z}_{1}),\bm{t}_{2}(\bm{z}_{2}))$.
@@
 To capture the non-linearity between input samples and their label distributions, deep neural networks can be leveraged to construct $\bm{t}_{2}(\cdot)$.
-The input can also be extended beyond the vector-format samples to data with special structures, like images, texts and graphs, where we can respectively leverage Convolutional Neural Networks (CNNs) (Venkatesan and Li, 2017), Transformers (Vaswani et al., 2017) and Graph Neural Networks (GNNs) (Kipf and Welling, 2017) to construct $\bm{t}_{2}(\cdot)$ for end-to-end learning.
-The conditional distribution formulation $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ also provides a closed form of mean, variance and co-variance of the label distribution probabilities conditioned on the input sample $\bm{x}$.
+The input can also be extended beyond the vector-format samples to data with special structures, like images, texts and graphs, where we can respectively leverage Convolutional Neural Networks (CNNs) [Venkatesan and Li, 2017], Transformers [Vaswani et al., 2017] and Graph Neural Networks (GNNs) [Kipf and Welling, 2017] to construct $\bm{t}_{2}(\cdot)$ for end-to-end learning.
+The conditional distribution formulation $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ also provides a closed form of mean, variance and co-variance of each label’s composition ratio in describing the conditioning sample $\bm{x}$.
 About this, we have the following theorem:
@@
 (6) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq.
-(10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
+(9), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the $r$th label’s composition ratio, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
 $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{F}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$ (10)
@@
 We can use the variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to quantify label distribution prediction uncertainties.
-We can also use $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ with $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to construct confidence intervals for label distribution predictions by applying Chebyshev’s inequality (Grimmett and Stirzaker, 2020).
+We can also use $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ with $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to construct confidence intervals for label distribution predictions by applying Chebyshev’s inequality [Grimmett and Stirzaker, 2020].
 The covariance $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$ is helpful for us to understand the correlations between two different labels.
-More importantly, all the statistics are conditioned on the input sample $\bm{x}$, guiding us to make instance-wise inferences.
+More importantly, all the statistics are conditioned on the given sample $\bm{x}$, guiding us to make instance-wise decisions.
 The conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ relies on the parameters $\bm{V}$ and $\bm{\Theta}$, as well as the neural network parameters for constructing $\bm{t}_{2}$ (we also use $\bm{t}_{2}$ to denote the parameters without confusion).
@@
 $\min_{\bm{V},\bm{\Theta},\bm{t}_{2}}-\sum_{\bm{x}^{\prime}\in\mathcal{X}}\log\frac{\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})}{d\bm{\ell}}\bigg{|}_{\bm{x}=\bm{x}^{\prime},\bm{\ell}=\bm{\ell}_{\bm{x}^{\prime}}}.$ (16)
-There are numerous metrics to measure the consistency between two label distributions, like Chebyshev distance, Kullback-Leibler divergence and Cosine coefficient (Geng, 2016).
+There are numerous metrics to measure the consistency between two label distributions, like Chebyshev distance, Kullback-Leibler divergence and Cosine coefficient [Geng, 2016].
 Instead of optimizing these metrics, the MLE based learning objective in Eq.
 (16) provides an alternative way to train the LDL model.
-The fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ can be applied to various downstream tasks, and its closed-form mean is able to offer accurate label distribution predictions that are generic to varying evaluation metrics.
-Algorithm Description and Time Complexity
+The fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ can be applied to various downstream tasks for quantifying the uncertainty of label distribution predictions.
+Algorithm Description and Time Complexity.
 We train the SNEFY-LDL model with stochastic gradient descent.
@@
 Taking the epoch number as a constant and assuming the latent layers of $\bm{t}_{2}$ have neurons in the same scale as the neuron number in the last layer $D_{2}$, the time complexity of Algorithm 1 is $\mathcal{O}(N(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2}))$, which is linear to the number of training samples $N$, making the algorithm able to scale to large datasets.
-For any sample $\bm{x}$ with an unknown label distribution, its label distribution can be predicted fast in time complexity $\mathcal{O}(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2})$ using the closed-form mean formulation in Eq.
+For any sample $\bm{x}$ with an unknown label distribution, its label distribution mean and variance can be computed in time complexity $\mathcal{O}(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2})$ using the closed-form formulations in Eq.
+(10) and Eq.
 (12).
@@
 10:  return optimized $(\bm{V},\bm{\Theta},\bm{t}_{2})$.
+Dataset #Examples #Features #Labels
+Movie 7,755 1,869 5
+Natural Scene 2,000 294 9
+SBU_3DFE 2,500 243 6
+SJAFFE 213 243 6
+Table 1: Summary of the four benchmark datasets.
 6 Experiments
-In this section, extensive experiments are carried out to evaluate SNEFY-LDL’s performance in label distribution prediction and prediction uncertainty quantification.
-Benchmark Datasets
-We use fourteen datasets (Geng, 2016) to benchmark our experiments, including the Movie dataset containing label distributions on five movie rating scales (from one to five), the Natural Scene dataset with label distributions constructed by inconsistent multi-label ranking on natural scene images, the facial expression datasets SBU_3DFE and SJAFFE with label distributions on six emotions, as well as the ten Yeast series datasets built from ten biological experiments on a budding yeast, where labels are defined by different time points and label distributions are constructed by different gene expression levels at varying time points.
-The statistics of the fourteen benchmark datasets are summarized in Table 1.
-Baseline Methods
-We compare the proposed SNEFY-LDL model with eight state-of-the-art LDL baselines:
-  • PT-SVM (Geng, 2016) transforms LDL into a multi-class classification problem and uses SVM as the classifier.
-  • AA-KNN (Geng, 2016) adapts KNN from the single-label classification setting to the LDL setting.
-  • CPNN (Geng et al., 2013) uses a neural network to model the feature-label joint distribution.
-  • SA-BFGS (Geng, 2016) applies the maximum entropy model to characterize label distributions.
-  • LDLF (Shen et al., 2017) employs the differentiable decision trees to model label distributions.
-  • DF-LDL (González et al., 2021a) learns label distributions in a decomposition-fusion manner.
-  • LDL-SCL (Jia et al., 2021) considers the label distribution correlations between similar samples.
-  • LDL-LRR (Jia et al., 2023a) preserves the label importance ranking in label distribution modeling.
-No. Dataset #Examples #Features #Labels
-1 Movie 7,755 1,869 5
-2 Natural Scene 2,000 294 9
-3 SBU_3DFE 2,500 243 6
-4 SJAFFE 213 243 6
-5 Yeast_alpha 2,465 24 18
-6 Yeast_cdc 2,465 24 15
-7 Yeast_cold 2,465 24 14
-8 Yeast_diau 2,465 24 7
-9 Yeast_dtt 2,465 24 6
-10 Yeast_elu 2,465 24 6
-11 Yeast_heat 2,465 24 4
-12 Yeast_spo 2,465 24 4
-13 Yeast_spo5 2,465 24 3
-14 Yeast_spoem 2,465 24 2
-Table 1: Summary of the fourteen benchmark datasets.
-Experimental Settings
-Following (Geng, 2016), we use ten-fold cross-validation to evaluate the label distribution prediction performance, by randomly splitting each dataset into ten chunks, and iteratively choosing a chunk as the test set and taking the remaining nine chunks as the training set.
-LDL models are then trained on the training set and evaluated on the test set for ten rounds.
-Averaged performance is used to compare SNEFY-LDL with baseline methods.
-The closed-form mean in Eq.
-(11) is used to make comparisons on the label distribution prediction task with the baselines which all return point estimates on the simplex.
-However, we note that SNEFY-LDL is a full conditional density model and its utility is broader than the baselines, as we illustrate in the active learning and ensemble learning experiments.
-When implementing SNEFY-LDL, $m$ and $n$ are respectively set to 32 and 64, $D_{2}$ is set as equal to $n$ and a one-layer neural network with ReLU activation is used to construct $\bm{t}_{2}$.
-The model is trained for 100 epochs with batch size 16.
-Weight clipping (Arjovsky et al., 2017) is used to control $\bm{W}_{1}>-1/2$ after each parameter update.
-For the eight baseline models, we use the implementations in the PyLDL library [^1] https://github.com/SpriteMisaka/PyLDL and keep the default hyperparameters.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 0.9332 $\bullet$ 0.9279 $\bullet$ 0.9244 $\bullet$ 0.9339 $\bullet$ 0.8596 $\bullet$ 0.9357 $\spadesuit$ 0.9036 $\bullet$ 0.9353 $\spadesuit$ 0.9354 -0.0002
-Natural_Scene 0.7404 $\spadesuit$ 0.7112 $\bullet$ 0.6249 $\bullet$ 0.7168 $\bullet$ 0.7304 $\spadesuit$ 0.7377 $\spadesuit$ 0.6716 $\bullet$ 0.7455 $\spadesuit$ 0.7524 +0.0068
-SBU_3DFE 0.9244 $\bullet$ 0.9215 $\bullet$ 0.9169 $\bullet$ 0.9383 $\bullet$ 0.9359 $\bullet$ 0.9386 $\bullet$ 0.9249 $\bullet$ 0.9284 $\bullet$ 0.9427 +0.0042
-SJAFFE 0.9336 $\bullet$ 0.9490 $\spadesuit$ 0.9302 $\bullet$ 0.9543 $\spadesuit$ 0.9593 $\circ$ 0.9646 $\circ$ 0.9604 $\circ$ 0.9635 $\circ$ 0.9507 -0.0140
-Yeast_alpha 0.9941 $\bullet$ 0.9938 $\bullet$ 0.9940 $\bullet$ 0.9946 $\circ$ 0.9943 $\spadesuit$ 0.9946 $\circ$ 0.9945 $\circ$ 0.9946 $\circ$ 0.9943 -0.0004
-Yeast_cdc 0.9928 $\bullet$ 0.9924 $\bullet$ 0.9927 $\bullet$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9929 -0.0004
-Yeast_cold 0.9879 $\bullet$ 0.9872 $\bullet$ 0.9875 $\bullet$ 0.9885 $\circ$ 0.9876 $\bullet$ 0.9884 $\spadesuit$ 0.9871 $\bullet$ 0.9885 $\circ$ 0.9883 -0.0002
-Yeast_diau 0.9870 $\bullet$ 0.9867 $\bullet$ 0.9860 $\bullet$ 0.9879 $\circ$ 0.9873 $\spadesuit$ 0.9879 $\circ$ 0.9849 $\bullet$ 0.9879 $\circ$ 0.9873 -0.0006
-Yeast_dtt 0.9934 $\bullet$ 0.9934 $\bullet$ 0.9938 $\spadesuit$ 0.9941 $\circ$ 0.9936 $\spadesuit$ 0.9939 $\spadesuit$ 0.9931 $\spadesuit$ 0.9941 $\circ$ 0.9939 -0.0002
-Yeast_elu 0.9933 $\bullet$ 0.9932 $\bullet$ 0.9936 $\spadesuit$ 0.9940 $\circ$ 0.9939 $\circ$ 0.9941 $\circ$ 0.9939 $\circ$ 0.9940 $\circ$ 0.9936 -0.0005
-Yeast_heat 0.9871 $\bullet$ 0.9867 $\bullet$ 0.9874 $\bullet$ 0.9880 $\circ$ 0.9877 $\spadesuit$ 0.9881 $\circ$ 0.9873 $\spadesuit$ 0.9880 $\circ$ 0.9878 -0.0003
-Yeast_spo 0.9765 $\spadesuit$ 0.9728 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9746 $\bullet$ 0.9756 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9766 -0.0004
-Yeast_spo5 0.9736 $\bullet$ 0.9711 $\bullet$ 0.9737 $\bullet$ 0.9742 $\spadesuit$ 0.9690 $\bullet$ 0.9698 $\bullet$ 0.9640 $\bullet$ 0.9742 $\spadesuit$ 0.9744 +0.0002
-Yeast_spoem 0.9783 $\bullet$ 0.9764 $\bullet$ 0.9789 $\spadesuit$ 0.9790 $\spadesuit$ 0.9750 $\bullet$ 0.9694 $\bullet$ 0.8925 $\bullet$ 0.9790 $\spadesuit$ 0.9792 +0.0001
-Table 2: The comparison of label distribution prediction performance measured by Cosine coefficient $\uparrow$.
-Evaluation Metrics
-Following (Geng, 2016), we adopt six metrics to evaluate label distribution prediction performance: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter).
-By denoting $\bm{\ell}$ and $\hat{\bm{\ell}}$ as ground-truth and predicted label distribution vectors respectively, these metrics are defined as follows:
-  • $\mathrm{Cheby}(\bm{\ell},\hat{\bm{\ell}})=\|\bm{\ell}-\hat{\bm{\ell}}\|_{\infty}\downarrow$.
-  • $\mathrm{Clark}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{2}\downarrow$.
-  • $\mathrm{Canb}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{1}\downarrow$.
-  • $\mathrm{KL}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\bm{\ell}^{y}\log\frac{\bm{\ell}^{y}}{\hat{\bm{\ell}}^{y}}\downarrow.$
-  • $\mathrm{Cos}(\bm{\ell},\hat{\bm{\ell}})=\frac{\bm{\ell}^{\top}\hat{\bm{\ell}}}{\|\bm{\ell}\|_{2}\|\hat{\bm{\ell}}\|_{2}}\uparrow.$
-  • $\mathrm{Inter}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\min(\bm{\ell}^{y},\hat{\bm{\ell}}^{y})\uparrow.$
-For each metric, $\uparrow$ ($\downarrow$) indicates that higher (lower) scores imply better label distribution prediction performance.
-Label Distribution Prediction
-Table 2 shows the label distribution prediction performance measured by Cosine coefficient, as the averaged scores on ten-fold cross-validation.
-On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline.
-We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level.
-The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline.
-From Table 2, we can find that the proposed SNEFY-LDL achieves very competitive label distribution prediction performance, compared with the state-of-the-art baselines.
-In the majority of cases, the performance of SNEFY-LDL is no worse than the compared baselines.
-On four datasets, SNEFY-LDL achieves the best Cosine coefficient scores.
-While SNEFY-LDL does not perform best on the remaining datasets, its performance gap with regard to the best performer is very marginal and even negligible.
-The performance scores measured by the remaining five metrics are provided in the Appendix, which show the same trend.
-The competitive label distribution prediction performance implies that the proposed SNEFY-LDL model is able to accurately model the distribution of label distribution vectors, offering an effective tool for label distribution prediction.
+We conduct extensive experiments on conformal prediction, active learning and ensemble learning to verify SNEFY-LDL’s ability in LDL uncertainty quantification.
+Class Id bin size $=2$ bin size $=4$ bin size $=8$
+Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL
+1 0.8577$\pm$0.0310 0.8747$\pm$0.0243 0.6987$\pm$0.2685 0.8524$\pm$0.0298 0.6476$\pm$0.2505 0.8335$\pm$0.0318
+2 0.8512$\pm$0.0333 0.8867$\pm$0.0300 0.4878$\pm$0.3210 0.8723$\pm$0.0361 0.4581$\pm$0.2952 0.8277$\pm$0.0472
+3 0.8921$\pm$0.0208 0.8924$\pm$0.0183 0.1644$\pm$0.3396 0.8104$\pm$0.0576 0.1395$\pm$0.2904 0.7537$\pm$0.0788
+4 0.8964$\pm$0.0176 0.8804$\pm$0.0256 0.5403$\pm$0.3666 0.8801$\pm$0.0257 0.4897$\pm$0.3246 0.8188$\pm$0.0412
+5 0.8492$\pm$0.0300 0.8690$\pm$0.0344 0.7561$\pm$0.2065 0.5276$\pm$0.3452 0.6676$\pm$0.1905 0.4883$\pm$0.3184
+6 0.8806$\pm$0.0248 0.8890$\pm$0.0230 0.2005$\pm$0.3537 0.8323$\pm$0.1615 0.1817$\pm$0.3193 0.8149$\pm$0.1582
+7 0.8447$\pm$0.0335 0.8662$\pm$0.0303 0.5131$\pm$0.3417 0.8499$\pm$0.0340 0.4827$\pm$0.3210 0.7697$\pm$0.0754
+8 0.8427$\pm$0.0329 0.8897$\pm$0.0218 0.5107$\pm$0.3399 0.8893$\pm$0.0219 0.4840$\pm$0.3175 0.8753$\pm$0.0245
+9 0.8107$\pm$0.0548 0.7681$\pm$0.0472 0.1443$\pm$0.2992 0.4513$\pm$0.2982 0.1242$\pm$0.2609 0.4096$\pm$0.2741
+Table 2: The conformal prediction performance measured by the FSC metric on the Natural_Scene dataset.
+Class Id bin size $=2$ bin size $=4$ bin size $=8$
+Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL
+1 0.8913$\pm$0.0205 0.8948$\pm$0.0175 0.8407$\pm$0.0511 0.8836$\pm$0.0189 0.5391$\pm$0.3022 0.8277$\pm$0.0675
+2 0.8816$\pm$0.0248 0.8658$\pm$0.0297 0.8703$\pm$0.0266 0.8472$\pm$0.0364 0.8475$\pm$0.0326 0.8210$\pm$0.0435
+3 0.8737$\pm$0.0278 0.8963$\pm$0.0175 0.8552$\pm$0.0292 0.8785$\pm$0.0258 0.8290$\pm$0.0349 0.8505$\pm$0.0297
+4 0.8848$\pm$0.0238 0.8879$\pm$0.0222 0.8274$\pm$0.0367 0.8725$\pm$0.0270 0.7193$\pm$0.1287 0.7261$\pm$0.1500
+5 0.8937$\pm$0.0177 0.8955$\pm$0.0184 0.8150$\pm$0.0462 0.8381$\pm$0.0403 0.0793$\pm$0.2388 0.4199$\pm$0.3376
+6 0.8911$\pm$0.0191 0.8920$\pm$0.0205 0.8625$\pm$0.0366 0.8334$\pm$0.0453 0.6025$\pm$0.2969 0.5048$\pm$0.3136
+Table 3: The conformal prediction performance measured by the FSC metric on the SBU_3DFE dataset.
+Benchmark Datasets.
+We use four datasets [Geng, 2016] to benchmark our experiments, including the Movie dataset containing label distributions on five movie rating scales, the Natural Scene dataset with label distributions constructed by inconsistent multi-label ranking on natural scene images, the facial expression datasets SBU_3DFE and SJAFFE with label distributions on six emotions.
+The statistics of the four benchmark datasets are summarized in Table 1.
+Implementation Details.
+When implementing SNEFY-LDL, $n$ and $m$ are respectively set to 64 and 32, $D_{2}$ is set as equal to $n$ and a one-layer neural network with ReLU activation is used to construct $\bm{t}_{2}$.
+The model is trained for 100 epochs with batch size 64 for conformal prediction and batch size 16 for active learning and ensemble learning.
+Weight clipping [Arjovsky et al., 2017] is used to control $\bm{W}_{1}>-1/2$ after each parameter update.
+6.1 Conformal Prediction
+As shown in Theorem 3, with the trained SNEFY-LDL model, given a new sample $\bm{x}$, we can get the closed-form conditional mean for the $r$th label’s composition ratio as $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ in Eq.
+(10) and the closed-form conditional variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ in Eq.
+(12).
+By applying the Chebyshev’s inequality [Grimmett and Stirzaker, 2020], we can construct a confidence interval for the $r$th label’s composition ratio in describing $\bm{x}$, $\bm{\ell}^{y_{r}}_{\bm{x}}$, which can be formally stated as
+$\mathrm{P}(|\bm{\ell}^{y_{r}}_{\bm{x}}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]|\leq k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]})\geq 1-\frac{1}{k^{2}},$ (17)
+i.e., the confidence interval of $\bm{\ell}^{y_{r}}_{\bm{x}}$ at the $1-1/k^{2}$ level is
+$\left[\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]-k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]},\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]+k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}\right],$
+which can be further calibrated as a conformal prediction task for the one-dimensional uncertainty estimate $\bm{\ell}^{y_{r}}_{\bm{x}}$.
+According to Angelopoulos and Bates [2021], on a calibration set with $N_{\text{cal}}$ samples, we can define a calibration score function as
+$s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})=\frac{\lvert\bm{\ell}^{y_{r}}_{\bm{x}}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]\rvert}{k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}},$ (18)
+where $\bm{\ell}^{y_{r}}_{\bm{x}}$ is the ground-truth value of the $r$th label’s composition ratio in describing the calibration sample $\bm{x}$.
+By computing the calibration scores of all calibration samples, we can get the $\lceil(1-1/k^{2})(N_{\text{cal}}+1)\rceil/N_{\text{cal}}$ quantile of the calibration scores as $\hat{q}_{y_{r}}$.
+For a new sample $\bm{x}$, the calibrated confidence interval for $\bm{\ell}^{y_{r}}_{\bm{x}}$ at the $1-1/k^{2}$ level is
+$\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})=\left[\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]-k\cdot\hat{q}_{y_{r}}\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]},\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]+k\cdot\hat{q}_{y_{r}}\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}\right].$ (19)
 Method Cheby $\downarrow$ Clark $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
-Random 0.1490$\pm$0.0098 $\bullet$ 0.6359$\pm$0.0247 $\bullet$ 1.1999$\pm$0.0448 $\bullet$ 0.1409$\pm$0.0136 $\bullet$ 0.9054$\pm$0.0093 $\bullet$ 0.7950$\pm$0.0089 $\bullet$
-Kmeans 0.1456$\pm$0.0076 $\bullet$ 0.6242$\pm$0.0176 $\bullet$ 1.1815$\pm$0.0320 $\bullet$ 0.1361$\pm$0.0072 $\bullet$ 0.9079$\pm$0.0068 $\bullet$ 0.7981$\pm$0.0077 $\bullet$
-CoreSet 0.1484$\pm$0.0164 $\bullet$ 0.6322$\pm$0.0383 $\bullet$ 1.1962$\pm$0.0720 $\bullet$ 0.1399$\pm$0.0179 $\bullet$ 0.9042$\pm$0.0158 $\bullet$ 0.7941$\pm$0.0171 $\bullet$
-Graph Density 0.1428$\pm$0.0059 $\bullet$ 0.6175$\pm$0.0197 $\bullet$ 1.1671$\pm$0.0346 $\bullet$ 0.1326$\pm$0.0068 $\bullet$ 0.9108$\pm$0.0057 $\bullet$ 0.8013$\pm$0.0076 $\bullet$
-Dirichlet 0.1472$\pm$0.0090 $\bullet$ 0.6282$\pm$0.0245 $\bullet$ 1.1899$\pm$0.0470 $\bullet$ 0.1376$\pm$0.0116 $\bullet$ 0.9064$\pm$0.0085 $\bullet$ 0.7959$\pm$0.0099 $\bullet$
-SNEFY 0.1350$\pm$0.0030 0.5981$\pm$0.0134 1.1283$\pm$0.0254 0.1209$\pm$0.0049 0.9191$\pm$0.0030 0.8098$\pm$0.0042
-Table 3: The label distribution active learning performance on the Movie dataset.
-Label Distribution Active Learning
-In addition, we conduct experiments on label distribution active learning to evaluate SNEFY-LDL’s performance in uncertainty quantification.
-We choose the Movie dataset, randomly select 400 labeled training samples to form the initial labeled pool and take the remaining training samples as the unlabeled pool on each fold of cross-validation.
-The SNEFY-LDL model is first trained on the initial labeled pool.
-We then use different active learning strategies to pick up 200 informative samples from the unlabeled pool and query their labels.
-After augmenting the initial labeled pool with the 200 queried samples, we re-train the SNEFY-LDL model and evaluate its label distribution prediction performance on the test sets.
+Random 0.1490$\pm$0.0098 0.6359$\pm$0.0247 1.1999$\pm$0.0448 0.1409$\pm$0.0136 0.9054$\pm$0.0093 0.7950$\pm$0.0089
+Kmeans 0.1456$\pm$0.0076 0.6242$\pm$0.0176 1.1815$\pm$0.0320 0.1361$\pm$0.0072 0.9079$\pm$0.0068 0.7981$\pm$0.0077
+CoreSet 0.1484$\pm$0.0164 0.6322$\pm$0.0383 1.1962$\pm$0.0720 0.1399$\pm$0.0179 0.9042$\pm$0.0158 0.7941$\pm$0.0171
+Graph Density 0.1428$\pm$0.0059 0.6175$\pm$0.0197 1.1671$\pm$0.0346 0.1326$\pm$0.0068 0.9108$\pm$0.0057 0.8013$\pm$0.0076
+Dirichlet 0.1472$\pm$0.0090 0.6282$\pm$0.0245 1.1899$\pm$0.0470 0.1376$\pm$0.0116 0.9064$\pm$0.0085 0.7959$\pm$0.0099
+SNEFY-LDL 0.1350$\pm$0.0030 0.5981$\pm$0.0134 1.1283$\pm$0.0254 0.1209$\pm$0.0049 0.9191$\pm$0.0030 0.8098$\pm$0.0042
+Table 4: The label distribution active learning performance on the Motive dataset.
+Method Cheby $\downarrow$ Clark $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
+Random 0.3591$\pm$0.0197 2.4871$\pm$0.0280 6.9131$\pm$0.1456 0.9835$\pm$0.0766 0.6554$\pm$0.0411 0.4528$\pm$0.0336
+Kmeans 0.3690$\pm$0.0163 2.4929$\pm$0.0200 6.9407$\pm$0.0808 1.0498$\pm$ 0.1078 0.6318$\pm$0.0375 0.4361$\pm$0.0228
+CoreSet 0.3629$\pm$0.0193 2.4842$\pm$0.0233 6.8882$\pm$0.1089 1.0102$\pm$0.1078 0.6442$\pm$0.0489 0.4485$\pm$0.0340
+Graph Density 0.3700$\pm$0.0242 2.4958$\pm$0.0277 6.9643$\pm$0.1394 1.0471$\pm$0.0991 0.6230$\pm$0.0453 0.4294$\pm$0.0321
+Dirichlet 0.3720$\pm$0.0213 2.4951$\pm$0.0226 6.9502$\pm$0.0977 1.0628$\pm$0.0909 0.6249$\pm$0.0336 0.4326$\pm$0.0192
+SNEFY-LDL 0.3474$\pm$0.0160 2.4807$\pm$0.0194 6.8755$\pm$0.0845 0.9244$\pm$0.0648 0.6819$\pm$0.0285 0.4718$\pm$0.0203
+Table 5: The label distribution active learning performance on the Natural_Scene dataset.
+As a baseline, we extend the competitive SA-BFGS [Geng, 2016] algorithm by modeling the distribution of (rather than point-estimating) the label distribution vectors.
+We model the distribution using the Dirichlet distribution centered at the point prediction.
+Following the same routine of conformal prediction, we can also construct calibrated confidence intervals for label affiliation probabilities given new samples.
+Following Angelopoulos and Bates [2021], we use the Feature-Stratified Coverage (FSC) metric to evaluate the adaptivity of the constructed confidence intervals in Eq.(19), which categorizes the test samples into different groups by dividing the numeric values at the first feature dimension into different bins, then compute the coverage rate of the confidence intervals in each group, and picks up the lowest group-level coverage rate as the final metric value.
+We select the Natural_Scene and SBU_3DFE datasets with a fair number of examples and labels to test the performance of conformal prediction.
+We randomly split each dataset into the training, calibration and test sets according to the ratio of 50%/25%/25% for 100 times and report the average FSC scores with bin size equal to 2, 4 and 8.
+In the experiment, we aim to construct 90% level confidence intervals by setting $0.11/k^{2}=$ 0.1, which means that FSC scores closer to 0.9 indicates better conformal prediction performance.
+Tables 2-3 compare the conformal prediction performance measured by the FSC metric on the Natural_Scene and SBU_3DFE datasets.
+For each comparison between SNEFY-LDL and Dirichlet, the better performer is highlighted by boldface.
+From the tables, we can find that the proposed SNEFY-LDL model outperforms the Dirichlet baseline in most cases.
+SNEFY-LDL constructs a multimodal distribution to model the distribution of label distribution vectors on the probability simplex space, which is more flexible than the unimodal Dirichlet, and contributes to the label distribution confidence intervals with greater adaptivity.
+6.2 Active Learning
+To further evaluate SNEFY-LDL’s performance in uncertainty quantification, we choose the Movie and Natural_Scene datasets to conduct the active learning experiments.
+We first randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times.
+For each training-test set split, we randomly select 400 labeled training samples to form the initial labeled pool and take the remaining training samples as the unlabeled pool.
+We first train a SNEFY-LDL model with the labeled samples in the initial labeled pool.
+We then use different active learning strategies to pick up 100 informative samples from the unlabeled pool and query their labels.
+After augmenting the initial labeled pool with the 100 queried samples, we re-train the SNEFY-LDL model and evaluate the performance of its label distribution predictions produced by the closed-form conditional mean in Eq.
+(10).
+Following Geng [2016], the label distribution prediction performance is evaluated by the following six metrics: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter).
+The average scores on the ten random training-test splits are reported.
 Six different active learning strategies are compared:
-  • Random (Zhan et al., 2022) randomly selects 200 samples from the unlabeled pool.
-  • Kmeans (Zhdanov, 2019) selects samples close to the cluster centroids generated by the Kmeans clustering (MacKay, 2003) in feature space.
-  • CoreSet (Sener and Savarese, 2018) selects the $k$-center samples (Har-Peled, 2011) as representative unlabeled samples, which is a variant of Kmeans.
-  • Graph Density (Ebert et al., 2012) selects highly connected samples in the constructed KNN graph (Preparata and Shamos, 2012).
-  • Dirichlet models the distribution of label distributions using a Dirichlet distribution (Ng et al., 2011) centered at predicted label distributions and selects samples with the largest differential entropies (Cover, 1999).
-  • SNEFY-LDL uses importance sampling (Kloek and Van Dijk, 1978) to estimate differential entropies of the conditional distributions modeled by SNEFY-LDL and picks up samples having the largest differential entropies.
-Among the six active learning strategies, Kmeans, CoreSet and Graph Density are representativeness based methods, which rely on only sample features by characterizing samples’ geometric properties in feature space, while Dirichlet and SNEFY-LDL are the uncertainty based methods we have contrived, which leverage the predictions of the initially trained SNEFY-LDL model.
-Table 3 shows the performance of different active learning strategies, where the best performer is highlighted by boldface and $\bullet$ indicates that the annotated performer is inferior to the best performer at 0.05 significance level.
-As is shown in Table 3, SNEFY-LDL consistently achieves the best performance in terms of all metrics.
-By accurately evaluating label distribution prediction uncertainties, SNEFY-LDL can pick up more informative unlabeled samples than the naive uncertainty quantification strategy, Dirichlet, as well as the representativeness based active learning strategies, Kmeans, CoreSet and Graph Density, which are even sometimes inferior to the Random strategy.
+  • Random [Zhan et al., 2022] randomly selects 100 samples from the unlabeled pool.
+  • Kmeans [Zhdanov, 2019] selects samples close to the cluster centroids generated by the Kmeans clustering [MacKay, 2003] in feature space.
+  • CoreSet [Sener and Savarese, 2018] selects the $k$-center samples [Har-Peled, 2011] as representative unlabeled samples, which is a variant of Kmeans.
+  • Graph Density [Ebert et al., 2012] selects highly connected samples in the constructed KNN graph [Preparata and Shamos, 2012].
+  • Dirichlet models the distribution of label distributions using a Dirichlet distribution [Ng et al., 2011] centered at predicted label distributions and selects samples with the largest differential entropy scores [Cover, 1999].
+  • SNEFY-LDL uses importance sampling [Kloek and Van Dijk, 1978] to estimate the differential entropy values of the conditional distributions modeled by SNEFY-LDL and picks up samples that have the largest differential entropy scores.
 Base Learner Bagging Cheby $\downarrow$ Clark $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
-SA-BFGS Average 0.1178$\pm$0.0020$\bullet$ 0.3743$\pm$0.0068$\bullet$ 0.7948$\pm$0.0163$\bullet$ 0.0641$\pm$0.0022$\bullet$ 0.9370$\pm$0.0020$\bullet$ 0.8575$\pm$0.0028$\bullet$
+SA-BFGS Average 0.1178$\pm$0.0020 0.3743$\pm$0.0068 0.7948$\pm$0.0163 0.0641$\pm$0.0022 0.9370$\pm$0.0020 0.8575$\pm$0.0028
 Weighted 0.1137$\pm$0.0023 0.3625$\pm$0.0062 0.7686$\pm$0.0149 0.0604$\pm$0.0021 0.9406$\pm$0.0020 0.8624$\pm$0.0026
-DF-LDL Average 0.1203$\pm$0.0017$\bullet$ 0.3762$\pm$0.0059$\bullet$ 0.8040$\pm$0.0146$\bullet$ 0.0657$\pm$0.0019$\bullet$ 0.9353$\pm$0.0018$\bullet$ 0.8557$\pm$0.0025$\bullet$
+DF-LDL Average 0.1203$\pm$0.0017 0.3762$\pm$0.0059 0.8040$\pm$0.0146 0.0657$\pm$0.0019 0.9353$\pm$0.0018 0.8557$\pm$0.0025
 Weighted 0.1152$\pm$0.0024 0.3617$\pm$0.0070 0.7715$\pm$0.0168 0.0609$\pm$0.0024 0.9399$\pm$0.0023 0.8617$\pm$0.0030
-LDL-SCL Average 0.1256$\pm$0.0020$\bullet$ 0.3828$\pm$0.0047$\bullet$ 0.8260$\pm$0.0118$\bullet$ 0.0699$\pm$0.0021$\bullet$ 0.9315$\pm$0.0018$\bullet$ 0.8519$\pm$0.0020$\bullet$
+LDL-SCL Average 0.1256$\pm$0.0020 0.3828$\pm$0.0047 0.8260$\pm$0.0118 0.0699$\pm$0.0021 0.9315$\pm$0.0018 0.8519$\pm$0.0020
 Weighted 0.1246$\pm$0.0023 0.3772$\pm$0.0048 0.8147$\pm$0.0114 0.0684$\pm$0.0022 0.9330$\pm$0.0019 0.8540$\pm$0.0020
-LDL-LRR Average 0.1269$\pm$0.0021$\bullet$ 0.3966$\pm$0.0052$\bullet$ 0.8478$\pm$0.0131$\bullet$ 0.0730$\pm$0.0022$\bullet$ 0.9285$\pm$0.0019$\bullet$ 0.8478$\pm$0.0023$\bullet$
+LDL-LRR Average 0.1269$\pm$0.0021 0.3966$\pm$0.0052 0.8478$\pm$0.0131 0.0730$\pm$0.0022 0.9285$\pm$0.0019 0.8478$\pm$0.0023
 Weighted 0.1250$\pm$0.0020 0.3916$\pm$0.0044 0.8373$\pm$0.0106 0.0710$\pm$0.0020 0.9305$\pm$0.0018 0.8498$\pm$0.0019
 Table 6: The label distribution ensemble learning performance on the SBU_3DFE dataset.
-Label Distribution Ensemble Learning
-We also conduct experiments on ensemble learning to further verify SNEFY-LDL’s ability in uncertainty quantification, by using Bagging (Breiman, 1996) as an exemplary ensemble learning paradigm.
-We choose the SBU_3DFE dataset, randomly select 50 samples from the training split of each fold of cross-validation for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of ensembled LDL models on the test sets.
-Four competitive LDL models are employed as base learners: SA-BFGS, DF-LDL, LDL-SCL and LDL-LRR, and two strategies are adopted to ensemble base learner predictions: 1) Average: aggregate the 25 base learner predictions with the uniform weight $1/25$, and 2) Weighted: weight each base learner prediction in proportion to its SNEFY-LDL conditional likelihood in an instance-wise manner.
-Table 4 compares the two different ensemble learning strategies, where the best strategy is highlighted by boldface and $\bullet$ indicates that the annotated strategy is inferior to the best strategy at 0.05 significance level.
-From Table 4, we can find that the Weighted strategy is significantly better than Average in terms of all metrics.
-This implies that SNEFY-LDL provides an effective mechanism to quantify the prediction reliability/uncertainties of base learners.
-Parameter Sensitivity Study
-We also conduct additional experiments to investigate SNEFY-LDL’s sensitivity with regard to its four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training.
-The experiments demonstrate that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges.
-Please check the Appendix for more details.
+Base Learner Bagging Cheby $\downarrow$ Clarky $\downarrow$ Canb $\downarrow$ KL $\downarrow$ Cos $\uparrow$ Inter $\uparrow$
+SA-BFGS Average 0.0889$\pm$0.0085 0.3180$\pm$0.0197 0.6529$\pm$0.0461 0.0406$\pm$0.0058 0.9613$\pm$0.0057 0.8890$\pm$0.0086
+Weighted 0.0842$\pm$0.0084 0.3118$\pm$0.0200 0.6390$\pm$0.0466 0.0385$\pm$0.0056 0.9636$\pm$0.0056 0.8923$\pm$0.0089
+DF-LDL Average 0.0951$\pm$0.0093 0.3385$\pm$0.0253 0.6958$\pm$0.0611 0.0456$\pm$0.0078 0.9566$\pm$0.0072 0.8818$\pm$0.0110
+Weighted 0.0881$\pm$0.0092 0.3189$\pm$0.0258 0.6525$\pm$0.0617 0.0408$\pm$0.0074 0.9612$\pm$0.0069 0.8895$\pm$0.0110
+LDL-SCL Average 0.0911$\pm$0.0090 0.3249$\pm$0.0219 0.6746$\pm$0.0493 0.0424$\pm$0.0063 0.9596$\pm$0.0062 0.8854$\pm$0.0092
+Weighted 0.0865$\pm$0.0095 0.3153$\pm$0.0231 0.6509$\pm$0.0540 0.0400$\pm$0.0064 0.9621$\pm$0.0064 0.8898$\pm$0.0100
+LDL-LRR Average 0.0888$\pm$0.0090 0.3189$\pm$0.0209 0.6538$\pm$0.0473 0.0408$\pm$0.0065 0.9612$\pm$0.0063 0.8890$\pm$0.0090
+Weighted 0.0846$\pm$0.0089 0.3124$\pm$0.0197 0.6392$\pm$0.0444 0.0387$\pm$0.0062 0.9634$\pm$0.0063 0.8922$\pm$0.0087
+Table 7: The label distribution ensemble learning performance on the SJAFFE dataset.
+Tables 4-5 compare the performance of different active learning strategies, where the best performer is highlighted by boldface.
+As is shown in the tables, SNEFY-LDL consistently achieves the best performance in terms of all metrics.
+By accurately evaluating label distribution prediction uncertainties, SNEFY-LDL can pick up more informative unlabeled samples than the naive uncertainty quantification strategy, Dirichlet, as well as the representativeness based active learning strategies, Kmeans, CoreSet and Graph Density, which are even sometimes inferior to the Random strategy.
+6.3 Ensemble Learning
+We also conduct experiments on ensemble learning to further verify SNEFY-LDL’s ability in uncertainty quantification, with the expectation that reliable base learners can be identified by the SNEFY-LDL probability modeling.
+Bagging [Breiman, 1996] is adopted as an exemplary ensemble learning paradigm.
+We choose the SEU_3DFE and SJAFFE datasets, and randomly split them into training and test sets according to the ratio of 90%/10%.
+For each training-test set split, we randomly select 50 samples from the training set for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of the ensembled LDL model on the test set.
+Four competitive LDL models are employed as base learners: SA-BFGS [Geng, 2016], DF-LDL [González et al., 2021a], LDL-SCL [Jia et al., 2021] and LDL-LRR [Jia et al., 2023a], and two strategies are adopted to ensemble base learner predictions: 1) Average: aggregate the 25 base learner predictions with the uniform weight $1/25$, and 2) Weighted: weight each base learner prediction in proportion to its corresponding SNEFY-LDL probability density conditioned on each test sample in an instance-wise manner.
+Tables 6-7 compare the two different ensemble learning strategies, where the best strategy is highlighted by boldface.
+From the tables, we find that the Weighted strategy is significantly better than Average in terms of all metrics.
+This implies that SNEFY-LDL provides an effective mechanism to quantify the reliability of base learners’ label distribution predictions so that the reliable base learners are highlighted to contribute to a better ensemble learning performance.
 7 Conclusion
-In this paper, we propose a novel LDL paradigm: estimate the distribution of label distribution vectors in the probability simplex, which brings a bird’s-eye view on the relative significance of all possible label distributions.
-Through uncovering the underlying relationship between SNEFY and LDL, we develop the SNEFY-LDL model that can provide a tractable formulation of the conditional distribution of label distribution vectors, enjoying great expressivity and high computational efficiency.
-SNEFY-LDL admits closed-form expressions for the distribution’s mean, variance and covariance that can be computed quickly, making SNEFY-LDL able to provide real-time responses to the downstream tasks and real-world applications.
-Experiments on label distribution prediction, as well as label distribution active learning and ensemble learning demonstrate the great utility of SNEFY-LDL for label distribution modeling.
+We propose a novel LDL paradigm: estimate the distribution of label distribution vectors on the probability simplex, which brings a bird’s-eye view on the relative significance of all possible label distributions.
+By uncovering the underlying relationship between SNEFY and LDL, we develop the SNEFY-LDL model that can provide a tractable formulation of the conditional distribution of label distribution vectors, enjoying great expressivity and high computational efficiency.
+SNEFY-LDL admits closed-form expressions for the distribution’s mean, variance and covariance, making SNEFY-LDL able to provide real-time responses in real-world applications.
+Experiments on conformal prediction, active learning and ensemble learning demonstrate the great utility of SNEFY-LDL for uncertainty-aware applications.
 Appendix A Theorem Proofs
@@
 As $\bm{W}_{1}>-1/2$ elementwise, $w_{1il}+w_{1jl}+1>0$.
-Assuming $\bm{\ell}$ follows a Dirichlet distribution (Ng, Tian, and Tang 2011) with parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{L})$, where $\alpha_{l}=w_{1il}+w_{1jl}+1>0$, its probability density, $\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})/{d\bm{\ell}}$, is in the form:
+Assuming $\bm{\ell}$ follows a Dirichlet distribution with parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{L})$, where $\alpha_{l}=w_{1il}+w_{1jl}+1>0$, its probability density, $\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})/{d\bm{\ell}}$, is in the form:
 $\frac{\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})}{d\bm{\ell}}=\frac{1}{\mathrm{B}(\bm{\alpha})}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1},$
 where $\mathrm{B}(\cdot)$ is the beta function. Considering the fact that $\int_{\Delta^{L-1}}\mathrm{P}_{\mathrm{Dir}}(d\bm{\ell})=1$,
-$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1}d\bm{\ell}=\mathrm{B}(\alpha).$
+$\int_{\Delta^{L-1}}\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{\alpha_{l}-1}d\bm{\ell}=\mathrm{B}(\bm{\alpha}).$
 That is to say
@@
 (6) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq.
-(10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
+(9), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the $r$th label’s composition ratio, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
 $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]=\frac{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x})\circ\bm{F}^{y_{r}})}{\mathrm{vec}(\bm{V}^{\top}\bm{V})^{\top}\mathrm{vec}(\bm{K}_{\bm{\Theta}}(\bm{x}))},$ (10)
@@
 Proof.
-For any fixed function $\varphi(\bm{\ell})$ of $\bm{\ell}$, its expectation with regard to the conditional SNEFY distribution in Eq. (7), $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]$, can be computed as
+For any fixed scalar function $\varphi(\bm{\ell})$ of $\bm{\ell}$ with $\varphi(\cdot):\Delta^{L-1}\rightarrow\mathbb{R}$, its expectation with regard to the conditional SNEFY distribution in Eq.
+(6), $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]$, can be computed as
 $\mathrm{E}[\varphi(\bm{\ell})|\bm{x}]=$ $\int_{\Delta^{L-1}}\varphi(\bm{\ell})\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$
@@
 […]
-$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}$
+$=\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\int_{\Delta^{L-1}}\varphi(\bm{\ell})\prod_{l=1}^{L}(\bm{\ell}^{y_{l}})^{w_{1il}+w_{1jl}}d\bm{\ell}.$
 As $\bm{W}_{1}>-1/2$ elementwise, $w_{1il}+w_{1jl}+1>0$.
@@
 Therefore,
-$\bm{\phi}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})$
-$\cdot$ $\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
+$\bm{\phi}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})=$ $\exp(\bm{w}_{2i}^{\top}\bm{t}_{2}(\bm{x})+\bm{w}_{2j}^{\top}\bm{t}_{2}(\bm{x})+b_{i}+b_{j})\cdot\frac{\prod_{l=1}^{L}\Gamma(1+w_{1il}+w_{1jl})}{\Gamma\big{(}L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})\big{)}}\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}]$
 $=$ $\bm{k}_{\bm{t}_{2}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})\mathrm{E}_{\mathrm{Dir}}[\varphi(\bm{\ell});\bm{w}_{1i},\bm{w}_{1j}].$
@@
 In particular, for $\varphi(\bm{\ell})=\bm{\ell}^{y_{r}}$, $\varphi(\bm{\ell})=(\bm{\ell}^{y_{r}})^{2}$, and $\varphi(\bm{\ell})=\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}}$ with $y_{r}\neq y_{s}$, we respectively have
-$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})}.$
-$\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$
+$\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{1+w_{1ir}+w_{1jr}}{L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})},$
+$\mathrm{E}_{\mathrm{Dir}}[(\bm{\ell}^{y_{r}})^{2};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(2+w_{1ir}+w_{1jr})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]},$
 $\mathrm{E}_{\mathrm{Dir}}[\bm{\ell}^{y_{r}}\cdot\bm{\ell}^{y_{s}};\bm{w}_{1i},\bm{w}_{1j}]=\frac{(1+w_{1ir}+w_{1jr})(1+w_{1is}+w_{1js})}{[L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})][1+L+\sum_{l=1}^{L}(w_{1il}+w_{1jl})]}.$
@@
 ∎
-Appendix B Label Distribution Prediction Results
-Tables 5-9 show the label distribution prediction performance measured by Chebyshev distance, Clark distance, Canberra metric, Kullback-Leibler divergence and Intersection, respectively, as the averaged scores on ten-fold cross-validation.
-On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline.
-We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level.
-The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline.
-The scores in Tables 5-9 consistently confirm that SNEFY-LDL is very competitive compared with the state-of-the-art baselines.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 0.1183 $\bullet$ 0.1193 $\bullet$ 0.1330 $\bullet$ 0.1163 $\spadesuit$ 0.1668 $\bullet$ 0.1135 $\circ$ 0.1444 $\bullet$ 0.1154 $\circ$ 0.1172 +0.0037
-Natural_Scene 0.3104 $\spadesuit$ 0.3130 $\spadesuit$ 0.3563 $\bullet$ 0.3194 $\spadesuit$ 0.3023 $\spadesuit$ 0.3077 $\spadesuit$ 0.3603 $\bullet$ 0.3046 $\spadesuit$ 0.3135 +0.0112
-SBU_3DFE 0.1300 $\bullet$ 0.1277 $\bullet$ 0.1394 $\bullet$ 0.1152 $\bullet$ 0.1197 $\bullet$ 0.1147 $\bullet$ 0.1303 $\bullet$ 0.1268 $\bullet$ 0.1110 -0.0037
-SJAFFE 0.1158 $\bullet$ 0.0975 $\spadesuit$ 0.1202 $\bullet$ 0.0921 $\spadesuit$ 0.0862 $\circ$ 0.0810 $\circ$ 0.0868 $\circ$ 0.0831 $\circ$ 0.0997 +0.0188
-Yeast_alpha 0.0140 $\bullet$ 0.0144 $\bullet$ 0.0140 $\bullet$ 0.0134 $\circ$ 0.0138 $\spadesuit$ 0.0134 $\circ$ 0.0135 $\circ$ 0.0134 $\circ$ 0.0137 +0.0003
-Yeast_cdc 0.0168 $\spadesuit$ 0.0172 $\bullet$ 0.0169 $\bullet$ 0.0162 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0162 $\circ$ 0.0167 +0.0005
-Yeast_cold 0.0529 $\bullet$ 0.0542 $\bullet$ 0.0538 $\bullet$ 0.0512 $\spadesuit$ 0.0535 $\bullet$ 0.0514 $\spadesuit$ 0.0543 $\bullet$ 0.0512 $\spadesuit$ 0.0518 +0.0006
-Yeast_diau 0.0386 $\spadesuit$ 0.0384 $\spadesuit$ 0.0407 $\bullet$ 0.0369 $\circ$ 0.0376 $\spadesuit$ 0.0369 $\circ$ 0.0405 $\bullet$ 0.0370 $\circ$ 0.0382 +0.0013
-Yeast_dtt 0.0381 $\bullet$ 0.0385 $\bullet$ 0.0370 $\spadesuit$ 0.0360 $\circ$ 0.0374 $\spadesuit$ 0.0365 $\spadesuit$ 0.0393 $\spadesuit$ 0.0360 $\circ$ 0.0369 +0.0009
-Yeast_elu 0.0171 $\bullet$ 0.0173 $\bullet$ 0.0169 $\spadesuit$ 0.0163 $\circ$ 0.0163 $\circ$ 0.0162 $\circ$ 0.0164 $\circ$ 0.0163 $\circ$ 0.0168 +0.0005
-Yeast_heat 0.0438 $\bullet$ 0.0440 $\bullet$ 0.0433 $\bullet$ 0.0422 $\spadesuit$ 0.0425 $\spadesuit$ 0.0420 $\circ$ 0.0435 $\spadesuit$ 0.0422 $\spadesuit$ 0.0425 +0.0006
-Yeast_spo 0.0589 $\bullet$ 0.0629 $\bullet$ 0.0593 $\bullet$ 0.0584 $\spadesuit$ 0.0612 $\bullet$ 0.0603 $\bullet$ 0.0592 $\bullet$ 0.0583 $\spadesuit$ 0.0583 -0.0000
-Yeast_spo5 0.0923 $\bullet$ 0.0952 $\bullet$ 0.0920 $\bullet$ 0.0913 $\spadesuit$ 0.0998 $\bullet$ 0.0973 $\bullet$ 0.1083 $\bullet$ 0.0912 $\spadesuit$ 0.0910 -0.0002
-Yeast_spoem 0.0884 $\bullet$ 0.0907 $\bullet$ 0.0874 $\spadesuit$ 0.0868 $\spadesuit$ 0.0932 $\bullet$ 0.1043 $\bullet$ 0.2315 $\bullet$ 0.0865 $\spadesuit$ 0.0865 -0.0000
-Table 5: The comparison of label distribution prediction performance measured by Chebyshev $\downarrow$.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 0.5293 $\circ$ 0.5312 $\spadesuit$ 0.5636 $\bullet$ 0.5251 $\circ$ 0.7024 $\bullet$ 0.5107 $\circ$ 0.6230 $\bullet$ 0.5214 $\circ$ 0.5370 +0.0263
-Natural_Scene 2.4441 $\circ$ 1.9084 $\circ$ 2.4878 $\bullet$ 2.4234 $\circ$ 2.4313 $\circ$ 2.4130 $\circ$ 2.3683 $\circ$ 2.4399 $\circ$ 2.4568 +0.5484
-SBU_3DFE 0.4067 $\bullet$ 0.4015 $\bullet$ 0.4151 $\bullet$ 0.3710 $\bullet$ 0.3770 $\bullet$ 0.3624 $\bullet$ 0.4025 $\bullet$ 0.3975 $\bullet$ 0.3429 -0.0195
-SJAFFE 0.4167 $\bullet$ 0.3477 $\spadesuit$ 0.4293 $\bullet$ 0.3589 $\spadesuit$ 0.3350 $\circ$ 0.3031 $\circ$ 0.3222 $\circ$ 0.3149 $\circ$ 0.3613 +0.0582
-Yeast_alpha 0.2218 $\bullet$ 0.2261 $\bullet$ 0.2239 $\bullet$ 0.2099 $\circ$ 0.2166 $\spadesuit$ 0.2094 $\circ$ 0.2122 $\circ$ 0.2099 $\circ$ 0.2182 +0.0088
-Yeast_cdc 0.2255 $\bullet$ 0.2317 $\bullet$ 0.2243 $\bullet$ 0.2158 $\circ$ 0.2173 $\circ$ 0.2157 $\circ$ 0.2179 $\circ$ 0.2159 $\circ$ 0.2219 +0.0062
-Yeast_cold 0.1443 $\bullet$ 0.1477 $\bullet$ 0.1458 $\bullet$ 0.1396 $\spadesuit$ 0.1456 $\bullet$ 0.1405 $\spadesuit$ 0.1479 $\bullet$ 0.1396 $\spadesuit$ 0.1411 +0.0015
-Yeast_diau 0.2096 $\bullet$ 0.2080 $\spadesuit$ 0.2178 $\bullet$ 0.2007 $\circ$ 0.2032 $\spadesuit$ 0.2001 $\circ$ 0.2213 $\bullet$ 0.2007 $\circ$ 0.2070 +0.0068
-Yeast_dtt 0.1041 $\bullet$ 0.1042 $\bullet$ 0.1008 $\spadesuit$ 0.0982 $\circ$ 0.1017 $\spadesuit$ 0.0994 $\spadesuit$ 0.1065 $\spadesuit$ 0.0982 $\circ$ 0.1006 +0.0025
-Yeast_elu 0.2114 $\bullet$ 0.2131 $\bullet$ 0.2073 $\spadesuit$ 0.1990 $\circ$ 0.2003 $\circ$ 0.1985 $\circ$ 0.2014 $\circ$ 0.1990 $\circ$ 0.2077 +0.0092
-Yeast_heat 0.1893 $\bullet$ 0.1908 $\bullet$ 0.1876 $\bullet$ 0.1826 $\spadesuit$ 0.1844 $\spadesuit$ 0.1815 $\circ$ 0.1882 $\bullet$ 0.1826 $\spadesuit$ 0.1839 +0.0024
-Yeast_spo 0.2530 $\bullet$ 0.2665 $\bullet$ 0.2533 $\bullet$ 0.2498 $\spadesuit$ 0.2594 $\bullet$ 0.2552 $\bullet$ 0.2530 $\bullet$ 0.2497 $\spadesuit$ 0.2494 -0.0003
-Yeast_spo5 0.1861 $\bullet$ 0.1915 $\bullet$ 0.1857 $\bullet$ 0.1840 $\spadesuit$ 0.1999 $\bullet$ 0.1951 $\bullet$ 0.2145 $\bullet$ 0.1838 $\spadesuit$ 0.1833 -0.0005
-Yeast_spoem 0.1313 $\bullet$ 0.1350 $\bullet$ 0.1299 $\spadesuit$ 0.1292 $\spadesuit$ 0.1392 $\bullet$ 0.1541 $\bullet$ 0.3628 $\bullet$ 0.1287 $\spadesuit$ 0.1288 +0.0001
-Table 6: The comparison of label distribution prediction performance measured by Clark distance $\downarrow$.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 1.0116 $\spadesuit$ 1.0205 $\spadesuit$ 1.0588 $\bullet$ 1.0055 $\spadesuit$ 1.4036 $\bullet$ 0.9796 $\circ$ 1.1795 $\bullet$ 0.9970 $\circ$ 1.0154 +0.0357
-Natural_Scene 6.7001 $\spadesuit$ 4.5465 $\circ$ 6.9670 $\bullet$ 6.6520 $\circ$ 6.6519 $\spadesuit$ 6.5907 $\circ$ 6.5539 $\circ$ 6.6785 $\circ$ 6.7202 +2.1737
-SBU_3DFE 0.8713 $\bullet$ 0.8306 $\bullet$ 0.9055 $\bullet$ 0.7827 $\bullet$ 0.8035 $\bullet$ 0.7638 $\bullet$ 0.8580 $\bullet$ 0.8485 $\bullet$ 0.7347 -0.0291
-SJAFFE 0.8604 $\bullet$ 0.7148 $\spadesuit$ 0.9001 $\bullet$ 0.7326 $\spadesuit$ 0.6820 $\circ$ 0.6188 $\circ$ 0.6648 $\circ$ 0.6421 $\circ$ 0.7473 +0.1285
-Yeast_alpha 0.7237 $\bullet$ 0.7388 $\bullet$ 0.7327 $\bullet$ 0.6818 $\circ$ 0.7051 $\spadesuit$ 0.6802 $\circ$ 0.6911 $\circ$ 0.6819 $\circ$ 0.7128 +0.0325
-Yeast_cdc 0.6794 $\bullet$ 0.7001 $\bullet$ 0.6733 $\spadesuit$ 0.6473 $\circ$ 0.6524 $\circ$ 0.6472 $\circ$ 0.6540 $\circ$ 0.6475 $\circ$ 0.6673 +0.0202
-Yeast_cold 0.2487 $\bullet$ 0.2550 $\bullet$ 0.2515 $\bullet$ 0.2403 $\spadesuit$ 0.2513 $\bullet$ 0.2424 $\spadesuit$ 0.2558 $\bullet$ 0.2403 $\spadesuit$ 0.2430 +0.0027
-Yeast_diau 0.4501 $\bullet$ 0.4476 $\spadesuit$ 0.4671 $\bullet$ 0.4308 $\circ$ 0.4367 $\spadesuit$ 0.4300 $\circ$ 0.4804 $\bullet$ 0.4308 $\circ$ 0.4449 +0.0149
-Yeast_dtt 0.1796 $\bullet$ 0.1789 $\bullet$ 0.1732 $\spadesuit$ 0.1689 $\circ$ 0.1750 $\spadesuit$ 0.1711 $\spadesuit$ 0.1832 $\spadesuit$ 0.1689 $\circ$ 0.1731 +0.0043
-Yeast_elu 0.6236 $\bullet$ 0.6281 $\bullet$ 0.6084 $\spadesuit$ 0.5831 $\circ$ 0.5876 $\circ$ 0.5820 $\circ$ 0.5905 $\circ$ 0.5832 $\circ$ 0.6116 +0.0296
-Yeast_heat 0.3780 $\bullet$ 0.3834 $\bullet$ 0.3739 $\bullet$ 0.3641 $\circ$ 0.3688 $\spadesuit$ 0.3626 $\circ$ 0.3763 $\bullet$ 0.3642 $\circ$ 0.3670 +0.0044
-Yeast_spo 0.5199 $\bullet$ 0.5489 $\bullet$ 0.5223 $\bullet$ 0.5131 $\spadesuit$ 0.5346 $\bullet$ 0.5245 $\bullet$ 0.5218 $\bullet$ 0.5129 $\spadesuit$ 0.5143 +0.0014
-Yeast_spo5 0.2856 $\bullet$ 0.2945 $\bullet$ 0.2847 $\bullet$ 0.2826 $\spadesuit$ 0.3082 $\bullet$ 0.3003 $\bullet$ 0.3344 $\bullet$ 0.2824 $\spadesuit$ 0.2820 -0.0004
-Yeast_spoem 0.1828 $\bullet$ 0.1878 $\bullet$ 0.1807 $\spadesuit$ 0.1797 $\spadesuit$ 0.1934 $\bullet$ 0.2149 $\bullet$ 0.4967 $\bullet$ 0.1791 $\spadesuit$ 0.1792 +0.0001
-Table 7: The comparison of label distribution prediction performance measured by Canberra metric $\downarrow$.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 0.1011 $\bullet$ 0.1093 $\bullet$ 0.1113 $\bullet$ 0.1002 $\bullet$ 0.2357 $\bullet$ 0.0975 $\spadesuit$ 0.1380 $\bullet$ 0.0981 $\spadesuit$ 0.0984 +0.0009
-Natural_Scene 0.7608 $\spadesuit$ 1.0936 $\bullet$ 1.0466 $\bullet$ 0.8277 $\spadesuit$ 0.7832 $\spadesuit$ 0.7608 $\spadesuit$ 1.0478 $\bullet$ 0.7445 $\spadesuit$ 0.7704 +0.0259
-SBU_3DFE 0.0774 $\bullet$ 0.0803 $\bullet$ 0.0860 $\bullet$ 0.0629 $\bullet$ 0.0651 $\bullet$ 0.0620 $\bullet$ 0.0776 $\bullet$ 0.0730 $\bullet$ 0.0578 -0.0042
-SJAFFE 0.0701 $\bullet$ 0.0534 $\spadesuit$ 0.0740 $\bullet$ 0.0505 $\spadesuit$ 0.0446 $\circ$ 0.0375 $\circ$ 0.0419 $\circ$ 0.0391 $\circ$ 0.0523 +0.0148
-Yeast_alpha 0.0060 $\bullet$ 0.0063 $\bullet$ 0.0061 $\bullet$ 0.0055 $\circ$ 0.0058 $\spadesuit$ 0.0055 $\circ$ 0.0056 $\circ$ 0.0055 $\circ$ 0.0059 +0.0004
-Yeast_cdc 0.0075 $\bullet$ 0.0079 $\bullet$ 0.0076 $\bullet$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0071 $\circ$ 0.0070 $\circ$ 0.0073 +0.0004
-Yeast_cold 0.0128 $\bullet$ 0.0136 $\bullet$ 0.0132 $\bullet$ 0.0122 $\circ$ 0.0132 $\bullet$ 0.0123 $\spadesuit$ 0.0137 $\bullet$ 0.0122 $\circ$ 0.0124 +0.0002
-Yeast_diau 0.0142 $\bullet$ 0.0145 $\spadesuit$ 0.0151 $\bullet$ 0.0131 $\circ$ 0.0138 $\spadesuit$ 0.0132 $\circ$ 0.0165 $\bullet$ 0.0131 $\circ$ 0.0138 +0.0007
-Yeast_dtt 0.0069 $\bullet$ 0.0070 $\bullet$ 0.0065 $\spadesuit$ 0.0063 $\circ$ 0.0068 $\spadesuit$ 0.0064 $\spadesuit$ 0.0072 $\spadesuit$ 0.0063 $\circ$ 0.0065 +0.0002
-Yeast_elu 0.0069 $\bullet$ 0.0071 $\bullet$ 0.0066 $\spadesuit$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0063 $\circ$ 0.0062 $\circ$ 0.0066 +0.0005
-Yeast_heat 0.0135 $\bullet$ 0.0139 $\bullet$ 0.0133 $\bullet$ 0.0126 $\circ$ 0.0130 $\spadesuit$ 0.0125 $\circ$ 0.0133 $\spadesuit$ 0.0126 $\circ$ 0.0129 +0.0004
-Yeast_spo 0.0251 $\spadesuit$ 0.0289 $\bullet$ 0.0258 $\bullet$ 0.0246 $\circ$ 0.0271 $\bullet$ 0.0259 $\bullet$ 0.0257 $\bullet$ 0.0246 $\circ$ 0.0251 +0.0005
-Yeast_spo5 0.0298 $\bullet$ 0.0326 $\bullet$ 0.0297 $\bullet$ 0.0292 $\spadesuit$ 0.0351 $\bullet$ 0.0338 $\bullet$ 0.0409 $\bullet$ 0.0292 $\spadesuit$ 0.0290 -0.0002
-Yeast_spoem 0.0253 $\bullet$ 0.0274 $\bullet$ 0.0247 $\spadesuit$ 0.0245 $\spadesuit$ 0.0301 $\bullet$ 0.0346 $\bullet$ 0.1475 $\bullet$ 0.0245 $\spadesuit$ 0.0243 -0.0002
-Table 8: The comparison of label distribution prediction performance measured by Kullback-Leibler divergence $\downarrow$.
-Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
-Movie 0.8324 $\spadesuit$ 0.8294 $\bullet$ 0.8196 $\bullet$ 0.8336 $\spadesuit$ 0.7572 $\bullet$ 0.8380 $\circ$ 0.7953 $\bullet$ 0.8353 $\circ$ 0.8333 -0.0048
-Natural_Scene 0.5549 $\spadesuit$ 0.5629 $\circ$ 0.4153 $\bullet$ 0.5491 $\spadesuit$ 0.5536 $\spadesuit$ 0.5566 $\spadesuit$ 0.5294 $\spadesuit$ 0.5595 $\circ$ 0.5343 -0.0286
-SBU_3DFE 0.8436 $\bullet$ 0.8484 $\bullet$ 0.8380 $\bullet$ 0.8600 $\bullet$ 0.8559 $\bullet$ 0.8624 $\bullet$ 0.8456 $\bullet$ 0.8477 $\bullet$ 0.8685 +0.0061
-SJAFFE 0.8523 $\bullet$ 0.8768 $\spadesuit$ 0.8466 $\bullet$ 0.8786 $\spadesuit$ 0.8865 $\circ$ 0.8959 $\circ$ 0.8876 $\circ$ 0.8928 $\circ$ 0.8729 -0.0229
-Yeast_alpha 0.9600 $\bullet$ 0.9592 $\bullet$ 0.9595 $\bullet$ 0.9624 $\circ$ 0.9611 $\spadesuit$ 0.9624 $\circ$ 0.9618 $\circ$ 0.9624 $\circ$ 0.9606 -0.0018
-Yeast_cdc 0.9552 $\bullet$ 0.9539 $\bullet$ 0.9557 $\spadesuit$ 0.9574 $\circ$ 0.9571 $\circ$ 0.9574 $\circ$ 0.9570 $\circ$ 0.9574 $\circ$ 0.9561 -0.0014
-Yeast_cold 0.9387 $\bullet$ 0.9371 $\bullet$ 0.9379 $\bullet$ 0.9408 $\spadesuit$ 0.9380 $\bullet$ 0.9403 $\spadesuit$ 0.9368 $\bullet$ 0.9408 $\spadesuit$ 0.9401 -0.0007
-Yeast_diau 0.9375 $\bullet$ 0.9379 $\spadesuit$ 0.9350 $\bullet$ 0.9403 $\circ$ 0.9394 $\spadesuit$ 0.9404 $\circ$ 0.9334 $\bullet$ 0.9403 $\circ$ 0.9383 -0.0021
-Yeast_dtt 0.9556 $\bullet$ 0.9559 $\bullet$ 0.9572 $\spadesuit$ 0.9583 $\circ$ 0.9568 $\spadesuit$ 0.9578 $\spadesuit$ 0.9547 $\spadesuit$ 0.9583 $\circ$ 0.9573 -0.0011
-Yeast_elu 0.9560 $\bullet$ 0.9557 $\bullet$ 0.9571 $\spadesuit$ 0.9589 $\circ$ 0.9585 $\circ$ 0.9589 $\circ$ 0.9583 $\circ$ 0.9589 $\circ$ 0.9568 -0.0021
-Yeast_heat 0.9379 $\bullet$ 0.9371 $\bullet$ 0.9386 $\bullet$ 0.9402 $\circ$ 0.9395 $\spadesuit$ 0.9405 $\circ$ 0.9382 $\bullet$ 0.9402 $\circ$ 0.9397 -0.0008
-Yeast_spo 0.9144 $\bullet$ 0.9094 $\bullet$ 0.9140 $\bullet$ 0.9155 $\spadesuit$ 0.9118 $\bullet$ 0.9135 $\bullet$ 0.9140 $\bullet$ 0.9156 $\spadesuit$ 0.9153 -0.0002
-Yeast_spo5 0.9077 $\bullet$ 0.9048 $\bullet$ 0.9080 $\bullet$ 0.9087 $\spadesuit$ 0.9002 $\bullet$ 0.9027 $\bullet$ 0.8917 $\bullet$ 0.9088 $\spadesuit$ 0.9090 +0.0002
-Yeast_spoem 0.9116 $\bullet$ 0.9093 $\bullet$ 0.9126 $\spadesuit$ 0.9132 $\spadesuit$ 0.9068 $\bullet$ 0.8957 $\bullet$ 0.7685 $\bullet$ 0.9135 $\spadesuit$ 0.9135 +0.0000
-Table 9: The comparison of label distribution prediction performance measured by Intersection $\uparrow$.
+Appendix B Experimental Details
+B.1 Conformal Prediction
+We select the Natural Scene and SBU_3DFE datasets for the conformal prediction experiments, as they have a relatively large number of labels that give more freedom in the change of label distribution compositions to warrant more uncertainties in label distribution predictions, and they also have enough samples to do the training/calibration/test set split.
+According to the ratio of 50%/25%/25%, we randomly split each dataset into the training, calibration and test sets.
+Here, we denote the training, calibration and test sets as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$, $\mathcal{S}^{\text{cal}}=\{(\bm{x}^{\text{cal}}_{i},\bm{\ell}_{\bm{x}^{\text{cal}}_{i}}):i=1,\cdots,N_{\text{cal}}\}$ and $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$ respectively, with $N_{\text{tr}}$, $N_{\text{cal}}$ and $N_{\text{te}}$ denoting the number of samples in the training, calibration and test sets respectively.
+For each split, a SNEFY-LDL is first trained on the training set $\mathcal{S}^{\text{tr}}$.
+The trained SNEFY-LDL is then leveraged to construct a 90% level confidence interval for each label’s composition ratio with SNEFY-LDL’s closed-form conditional mean and variance.
+The confidence intervals constructed by the trained SNEFY-LDL model are then calibrated with the calibration set.
+The adaptivity of the calibrated confidence intervals are finally evaluated on the test set, as an indicator of SNEFY-LDL’s ability in quantifying the prediction uncertainty of each label’s composition ratio.
+On the test set, the adaptivity of the calibrated confidence intervals $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ in Eq.
+(19) for the $r$th label’s composition ratio $\bm{\ell}^{y_{r}}$ is measured by the Feature-Stratified Coverage (FSC) metric [Angelopoulos and Bates, 2021].
+For calculating the FSC metric, we first bin the first feature of $\bm{x}$ for all $\bm{x}\in\mathcal{S}_{\text{te}}$ into a number of categories, $1,\cdots,G$, then categorize test samples in $\mathcal{S}_{\text{te}}$ into different groups $\{\mathcal{S}^{\text{te}}_{g}:g=1,\cdots,G\}$ according to the first feature’s category values.
+Here, $G$ is termed as bin size or group number.
+The FSC metric for the confidence intervals of the $r$th label’s composition ratio $\bm{\ell}^{y_{r}}$ is evaluated as the minimal coverage rate among the groups $\{\mathcal{S}^{\text{te}}_{g}:g=1,\cdots,G\}$:
+$\text{FSC}(\bm{\ell}^{y_{r}}):=\min_{g\in\{1,\cdots,G\}}\frac{1}{|\mathcal{S}_{g}^{\text{te}}|}\sum_{\bm{x}\in\mathcal{S}_{g}^{\text{te}}}\mathbbm{1}\left\{\bm{\ell}_{\bm{x}}^{y_{r}}\in\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})\right\}.$ (20)
+The detailed procedure of conformal prediction with SNEFY-LDL is described by Algorithm A1.
+Algorithm A1 Conformal Prediction with SNEFY-LDL
+Input: Training set $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$, calibration set $\mathcal{S}^{\text{cal}}=\{(\bm{x}^{\text{cal}}_{i},\bm{\ell}_{\bm{x}^{\text{cal}}_{i}}):i=1,\cdots,N_{\text{cal}}\}$ and test set $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, as well as the given label set $\mathcal{Y}=\{y_{1},y_{2},\cdots,y_{L}\}$.
+Parameter: $0.11/k^{2}=$ 0.1 for constructing 90% level confidence intervals.
+Output: FSC scores on the test set.
+1:  Train a SNEFY-LDL model on the training set $\mathcal{S}^{\text{tr}}$ with Algorithm 1;
+2:  for each sample $\bm{x}\in\mathcal{S}^{\text{cal}}$ do
+3:     for each label $y_{r}\in\mathcal{Y}$ do
+4:        Calculate the calibration score $s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})$ with Eq.(18) and the trained SNEFY-LDL model;
+5:     end for
+6:  end for
+7:  for each label $y_{r}\in\mathcal{Y}$ do
+8:     Calculate the $\lceil(1-1/k^{2})(N_{\text{cal}}+1)\rceil/N_{\text{cal}}$ quantile as $\hat{q}_{y_{r}}$ among the $N_{\text{cal}}$ calibration scores $s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})$ with $\bm{x}\in\mathcal{S}^{\text{cal}}$;
+9:  end for
+10:  for each sample $\bm{x}\in\mathcal{S}^{\text{te}}$ do
+11:     for each label $y_{r}\in\mathcal{Y}$ do
+12:        Construct the calibrated confidence interval $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ with Eq. (19) and the calculated quantile $\hat{q}_{y_{r}}$;
+13:     end for
+14:  end for
+15:  for each label $y_{r}\in\mathcal{Y}$ do
+16:     Evaluate the $\text{FSC}(\bm{\ell}^{y_{r}})$ score with Eq. (20) and the $N_{\text{te}}$ confidence intervals $\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})$ with $\bm{x}\in\mathcal{S}^{\text{te}}$;
+17:  end for
+18:  return the evaluated $\text{FSC}(\bm{\ell}^{y_{r}})$ scores for all $y_{r}\in\mathcal{Y}$.
+B.2 Active Learning
+The Motive and Natural Scene datasets are selected for active learning, as they are relatively sensitive to the label sparsity issue, more suitable to benchmark the performance change with informative samples labeled and augmented to the training data.
+We randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times.
+We denote the training set as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set as $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, where $N_{\text{tr}}$ and $N_{\text{te}}$ are respectively the number of samples in the training and test sets.
+For each training-test set split, we randomly select 400 labeled samples from the training set $\mathcal{S}^{\text{tr}}$ to form the initial labeled pool and take the remaining samples in $\mathcal{S}^{\text{tr}}$ as unlabeled samples.
+We first train a SNEFY-LDL model with the initial labeled pool.
+To achieve active learning with the trained SNEFY-LDL model, we first evaluate the differential entropy $H(\bm{x};\bm{V},\bm{\Theta})$ for each unlabeled sample $\bm{x}$ as
+$H(\bm{x};\bm{V},\bm{\Theta})=-\int_{\Delta^{L-1}}\left\{\log\frac{\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})}{d\bm{\ell}}\right\}\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta}),$ (21)
+then select 100 most informative unlabeled samples with the largest differential entropy values.
+After querying the labels of the 100 selected samples, we augment them into the initial labeled pool, and re-train another SNEFY-LDL model with the augmented labeled pool.
+With the re-trained SNEFY-LDL model, we can predict the label distribution vectors for samples in the test set given their feature vectors $\bm{x}$ by directly using the closed-form conditional mean in Eq.(10).
+As a criterion of active learning, the label distribution prediction performance on the test set is evaluated using the six metrics [Geng, 2016]: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter).
+Given the ground-truth and predicted label distribution vectors as $\bm{\ell}$ and $\hat{\bm{\ell}}$ respectively, the evaluation metrics are defined as follows:
+$\text{Cheby}(\bm{\ell},\hat{\bm{\ell}})=\|\bm{\ell}-\hat{\bm{\ell}}\|_{\infty}\downarrow,\quad\text{Clark}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{2}\downarrow,\quad\text{Canb}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{1}\downarrow,$ (22)
+$\text{KL}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\bm{\ell}^{y}\log\frac{\bm{\ell}^{y}}{\hat{\bm{\ell}}^{y}}\downarrow,\;\;\;\text{Cos}(\bm{\ell},\hat{\bm{\ell}})=\frac{\bm{\ell}^{\top}\hat{\bm{\ell}}}{\|\bm{\ell}\|_{2}\|\hat{\bm{\ell}}\|_{2}}\uparrow,\quad\;\;\text{Inter}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\min(\bm{\ell}^{y},\hat{\bm{\ell}}^{y})\uparrow.$
+For each metric, $\uparrow$ ($\downarrow$) indicates that higher (lower) scores imply better label distribution prediction performance.
+However, the differential entropy $H(\bm{x};\bm{V},\bm{\Theta})$ in Eq.
+(21) cannot be computed in a closed form.
+To overcome this difficulty, we adopt importance sampling [Kloek and Van Dijk, 1978] to approximately estimate the differential entropy values, where the uniform distribution is selected as the proposal distribution.
+The detailed procedure is provided in Algorithm A2.
+Algorithm A2 Differential Entropy Estimation with Importance Sampling
+Input: The feature vector of an unlabeled sample $\bm{x}$ and the conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ modeled by the trained SNEFY-LDL model, and the label set size $L$.
+Parameter: The number of sampling iterations $N_{\text{iter}}=1,000$.
+Output: The estimated differential entropy for the unlabeled sample $\bm{x}$, $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$.
+1:  for each iteration $i\in\{1,\cdots,N_{\text{iter}}\}$ do
+2:     Sample a label distribution vector ~$\tilde{\bm{\ell}}^{(i)}$ from the uniform distribution over the simplex $\Delta^{L-1}$ with probability density $q(\tilde{\bm{\ell}}^{(i)})=(L-1)!$;
+3:     Evaluate the probability density value $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})/d\bm{\ell}$ at =~$\bm{\ell}=\tilde{\bm{\ell}}^{(i)}$ as $p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})$ with Eq.
+(6) and the trained SNEFY-LDL model;
+4:  end for
+5:  Calculate the approximated differential entropy $\hat{H}(\bm{x};\bm{V},\bm{\Theta})=-\frac{1}{N_{\text{iter}}}\sum_{i=1}^{N_{\text{iter}}}\frac{p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})}{q(\tilde{\bm{\ell}}^{(i)})}\log p(\tilde{\bm{\ell}}^{(i)}|\bm{x};\bm{V},\bm{\Theta})$;
+6:  return the estimated differential entropy $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$ for the given unlabeled sample $\bm{x}$.
+The detailed procedure for active learning with SNEFY-LDL is described by Algorithm A3.
+B.3 Ensemble Learning
+The SBU_3DFE and SJAFFE datasets are selected for the ensemble learning experiments, as they have a relatively small number of features, more efficient to train a number of base learners and do ensemble prediction.
+We randomly split the two datasets into the training and test sets according to the ratio of 90%/10% for ten times.
+We denote the training set as $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set as $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$, where $N_{\text{tr}}$ and $N_{\text{te}}$ are respectively the number of samples in the training and test sets.
+For each training-test set split, we randomly select 50 samples from the training set for 25 rounds, train 25 base LDL learners with the selected samples, and evaluate the label distribution prediction performance of the ensembled LDL model on the test set.
+Four competitive LDL algorithms are employed to train base learners: SA-BFGS [Geng, 2016], DF-LDL [González et al., 2021a], LDL-SCL [Jia et al., 2021] and LDL-LRR [Jia et al., 2023a].
+To achieve ensemble learning with SNEFY-LDL, we first train a SNEFY-LDL model with the training set $\mathcal{S}^{\text{tr}}$, and then do the ensemble prediction by weighting the base learners according to the SNEFY-LDL conditional probability densities measured at their predictions given each test sample $\bm{x}\in\mathcal{S}^{\text{te}}$.
+The detailed procedure is shown in Algorithm A4.
+Algorithm A3 Active Learning with SNEFY-LDL
+Input: Training set $\mathcal{S}^{\text{tr}}=\{(\bm{x}^{\text{tr}}_{i},\bm{\ell}_{\bm{x}^{\text{tr}}_{i}}):i=1,\cdots,N_{\text{tr}}\}$ and test set $\mathcal{S}^{\text{te}}=\{(\bm{x}^{\text{te}}_{i},\bm{\ell}_{\bm{x}^{\text{te}}_{i}}):i=1,\cdots,N_{\text{te}}\}$.
+Parameters: The size of the initial labeled pool $N_{\text{initial}}=400$ and the number of queried samples $N_{\text{query}}=100$.
+Output: The label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter on the test set $\mathcal{S}^{\text{te}}$.
+1:  Randomly select $N_{\text{initial}}$ samples from the training set $\mathcal{S}^{\text{tr}}$ to form the labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$ and use the remaining samples to form the unlabeled pool $\mathcal{S}^{\text{tr}}_{\text{unlabel}}$;
+2:  Train a SNEFY-LDL model with the initial labeled pool;
+3:  for each sample $\bm{x}\in\mathcal{S}^{\text{tr}}_{\text{unlabel}}$ do
+4:     Estimate the differential entropy value $\hat{H}(\bm{x};\bm{V},\bm{\Theta})$ for $\bm{x}$ with Algorithm A2 and the trained SNEFY-LDL model;
+5:  end for
+6:  Select $N_{\text{query}}$ samples from $\mathcal{S}^{\text{tr}}_{\text{unlabel}}$ with the top-$N_{\text{query}}$ differential entropy values and query their label distributions;
+7:  Augment the $N_{\text{query}}$ queried samples into the labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$;
+8:  Re-train another SNEFY-LDL model with the augmented labeled pool $\mathcal{S}^{\text{tr}}_{\text{label}}$;
+9:  Evaluate the label distribution prediction performance of the re-trained SNEFY-LDL model on the test set $\mathcal{S}^{\text{te}}$ (as an average over all test samples) with the metrics of Cheby, Clark, Canb, KL, Cos and Inter defined in Eq.
+(22);
+10:  return the label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter.
+Algorithm A4 Ensemble Learning with SNEFY-LDL
+[…]
+Parameters: The number of samples for training base learners $N_{\text{sample}}=50$ and the number of base learners $N_{\text{base}}=25$.
+Output: The label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter on the test set $\mathcal{S}^{\text{te}}$.
+1:  for each iteration $i\in\{1,\cdots,N_{\text{base}}\}$ do
+2:     Randomly select $N_{\text{sample}}$ samples from the training set $\mathcal{S}^{\text{tr}}$;
+3:     Train a LDL base learner $B_{i}$ with the selected samples;
+4:  end for
+5:  Train a SNEFY-LDL model with the training set $\mathcal{S}^{\text{tr}}$;
+6:  for each sample $\bm{x}\in\mathcal{S}^{\text{te}}$ do
+7:     for each iteration $i\in\{1,\cdots,N_{\text{base}}\}$ do
+8:        Predict $\bm{x}$’s label distribution vector with base learner $B_{i}$ as ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$;
+9:        Evaluate the probability density value $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})/d\bm{\ell}$ at =~$\bm{\ell}=\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$ as $p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})$ with Eq.
+(6) and the trained SNEFY-LDL model;
+10:     end for
+11:     Predict $\bm{x}$’s label distribution vector ~$\tilde{\bm{\ell}}_{\bm{x}}$ as the weighted average of ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$, i.e., ~$\tilde{\bm{\ell}}_{\bm{x}}=\frac{\sum_{i=1}^{N_{\text{base}}}p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})\tilde{\bm{\ell}}^{(i)}_{\bm{x}}}{\sum_{i=1}^{N_{\text{base}}}p(\tilde{\bm{\ell}}^{(i)}_{\bm{x}}|\bm{x};\bm{V},\bm{\Theta})}$;
+12:  end for
+13:  Evaluate the performance of the ensembled label distribution predictions ~$\tilde{\bm{\ell}}^{(i)}_{\bm{x}}$ on the test set $\mathcal{S}^{\text{te}}$ (as an average over all test samples) with the metrics of Cheby, Clark, Canb, KL, Cos and Inter defined in Eq.
+(22);
+14:  return the label distribution prediction performance scores measured by Cheby, Clark, Canb, KL, Cos and Inter.
 Appendix C Parameter Sensitivity Study
-On the Movie and SBU_3DFE datasets, we take turns to study the sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training, by varying the studied hyperparameter in a predefined range and fixing the remaining three as default values at each turn.
-Figure 1 plots the performance change of SNEFY-LDL measured by Cosine coefficient with the varying values of the four hyperparameters.
-From Figure 1, we can find that the performance of SNEFY-LDL remains relatively stable when the hyperparameters vary in proper ranges.
+By choosing the Natural Scene dataset and the conformal prediction task, we take turns to study the sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$ and $m$, as well as the batch size and epoch number used for training, by varying the studied hyperparameter in a predefined range and fixing the remaining three as default values at each turn.
+For the conformal prediction task, the default values of $n$, $m$, batch size and epoch number are 64, 32, 64 and 100 respectively.
+Figure A1 plots the conformal prediction performance change of SNEFY-LDL measured by FSC with bin size equal to 2 and 4 when the values of hyperparameters vary in a range.
+From Figure A1, we can find that the performance of SNEFY-LDL remains relatively stable with the change of the four hyperparameters in most cases, except for the cases with $m=48$ and $96$ in Figure A1(b), as well as batch size $=32$ and $96$ in Figure A1(c), where the conformal prediction performance of SNEFY-LDL goes through a obvious drop.
 (a) $n$
@@
 (d) epoch number
-Figure 1: Parameter sensitivity.
+Figure A1: The sensitivity of SNEFY-LDL with regard to the four hyperparameters: $n$, $m$, batch size and epoch number.